### v3

Abstract
We develop an information-theoretic framework for bounding the supremum of stochastic processes, offering a simpler and sharper alternative to classical chaining and slicing arguments for generalization bounds. The key idea is a lifting argument that produces information-theoretic analogues of empirical process bounds, such as Dudley’s entropy integral. Lifting introduces permutation symmetry, yielding sharp bounds when the classical Dudley integral is loose. This gives a simple proof of the majorizing measure theorem via the sharpness of Dudley’s entropy integral for stationary processes, a result known well before the proof of the majorizing measure theorem. Furthermore, the information-theoretic formulation provides soft versions of classical localized complexity bounds in generalization theory, but is simpler and does not require the slicing argument. We apply this approach to empirical risk minimization over Sobolev ellipsoids, obtaining sharp convergence rates in settings where previous methods are suboptimal.
1 Introduction
Sharp control of empirical processes is central to modern statistical learning theory, underpinning generalization bounds, minimax rates, and risk analysis for high-dimensional estimators. Classical tools such as Dudley’s entropy integral have played a central role in this analysis, with applications ranging from nonparametric regression [41, 44] to sparse high-dimensional models [35]. Although certain special cases can be handled by simpler techniques, the localized version of Dudley’s integral [41] has remained a fundamental tool in many general settings. However, its application to statistical learning may still be limited since (i) Dudley’s integral, while widely used, can be suboptimal even for basic examples such as Sobolev ellipsoids, and (ii) the slicing (or peeling) technique for localization often requires delicate covering number estimates, sometimes invoking deep results from approximation theory or the local theory of Banach spaces, e.g. in sparse regression over the $\ell_{q}$ ball [35].
In this paper, we introduce a “lifting” method for deriving new information-theoretic generalization bounds, which overcomes the limitations discussed above. The key idea is to introduce replicated copies of a system and apply the method of types from information theory, so that traditional “hard” inequalities for the supremum of a stochastic process for an index set yield “soft” information-theoretic inequalities for a distribution on that set. The strength of the lifting approach lies in leveraging geometric insights involving covering and packing numbers, such as the chaining argument in Dudley’s integral, yielding new information-theoretic inequalities that are not easily derived from traditional information-theoretic methods. This method yields sharper localized generalization bounds and new results for empirical risk minimization in some settings.
For example, consider a Gaussian process $(X_{t})_{t\in T}$ and a measure $\mu$ on $T$. With the lifting approach, both Dudley’s integral and the majorizing measure theorem yield the following equivalences for Gaussian processes (up to universal constant factors):
$\sup\mathbb{E}[X_{Z}]\asymp\int_{0}^{\infty}\sqrt{\inf_{P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)}I(Z;\hat{Z})}d\sigma\asymp\int_{0}^{\infty}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (1)
where $\Pi_{\sigma}(\mu)$ denotes the set of couplings of $\mu$ and itself with mean squared error $\sigma^{2}$, I(Z;Z^)I($\hat{Z}$) is the mutual information, and $R_{\mu}(\sigma^{2})$ denotes the rate-distortion function (Definition 3). The supremum in (1) is over all conditional distributions of the index $Z$ given the process such that $P_{Z}=\mu$. While the analogy between the rate-distortion function $R_{\mu}(\sigma^{2})$ and covering numbers is well-known [12, 13], we are not aware of any prior literature explicitly stating (1). Previously, similar lifting arguments were used for problems in selected problems from network information theory [46, 26, 25, 27], but the application to empirical processes appears to be new. More generally, “soft” versions of traditional measure concentration techniques are known to provide sharper convergence in some hypothesis testing and information theory problems [23, 30, 24, 28].
We use information-theoretic inequalities such as (1) to control the generalization error in empirical risk minimization (ERM), and this approach offers several advantages over the traditional approach based on slicing and Dudley’s integral:
First, our approach yields sharper convergence rate in some basic estimation problems. The reason is that (1) is a sharp two-sided inequality for any $\mu$, which is perhaps surprising, since (1) is derived from the Dudley integral, and the latter is not sharp for some $T$. The reason for the sharpness of (1) is that the lifting argument transforms a general Gaussian process into a stationary one (i.e., invariant under the transitive group of permutations; see Definition 1), even if the original process lacks this symmetry. Since Dudley’s integral is sharp for stationary processes (see Section 2.1 for background), (1) inherits the sharpness, even if the original process is nonstationary. The sharpness of (1) also has an interesting consequence: by essentially maximizing over $\mu$ and applying the date processing inequality, we obtain a concise new proof of the nontrivial lower bound part of the classical majorizing measure theorem. This new proof of the majorizing measure theorem is conceptually simple, as it closely mirrors the argument used for the Dudley integral upper bound, essentially reversing its key steps. It may also be interpreted as saying that Fernique’s result about sharpness of Dudley’s inequality for stationary processes [17] already contains, in essence, Talagrand’s majorizing measure theorem [37], once combined with relatively standard information-theoretic techniques. Figure 1 illustrates the implication relations of these results.
Dudley’s IntegralMajorizing Measure TheoremRate-Distortion Integralstationary processliftingliftingdata processing
Figure 1: Arrows indicate implications of results, and the techniques/assumptions for the implications. For Gaussian processes, the majorizing measure theorem and the proposed rate-distortion integral are two-sided bounds, whereas Dudley’s integral is only an upper bound unless stationarity is assumed.
Second, our approach makes localization of generalization bounds more straightforward. Localization refers to the phenomenon where the index $Z$ of the process concentrates on a subset much smaller than the full index set $T$, so that taking the supremum over all of $T$ may significantly overestimate the typical value of the process at $Z$. For example, in empirical risk minimization (ERM), $Z$ denotes the estimator, which tends to concentrate near zero as the sample size increases. Traditionally, localization is achieved by the “slicing” (or “peeling”) argument, which yields sharp convergence rate for some empirical risk minimization problems [44]. Slicing typically involves controlling the modulus of continuity of the process through careful controls of the deviation magnitude for different slices. In contrast, using properties of the rate-distortion function, we can often directly upper bound the right side of (1) using the variance of $Z\sim\mu$, which is much smaller than the size of $T$. Similarly, tail probability bounds on $X_{Z}$, based on the moment generating function of $\|Z\|_{2}$, can also be derived using a lifting technique. This new method of localization does not require slicing.
To demonstrate our approach, we obtain new results for mean estimation via ERM when the parameter set is the image of the Sobolev ellipsoid under a Lipschitz map. Sobolev ellipsoid play a fundamental role in nonparametric statistics [40], whereas the Lipschitz map may represent a nonlinear feature map and is a technically challenging ingredient that necessitates the majorizing measure theorem machinery (see sub-Gaussian comparison (10)). Assuming smoothness parameter $\beta>1/2$, the traditional slicing method based on Dudley’s integral [41] yields an upper bound of order $\tfrac{1}{2\beta-1}n^{-2\beta/(2\beta+1)}$ for the mean squared error. In contrast, our method achieves the sharper rate $\tfrac{1}{\sqrt{2\beta-1}}n^{-2\beta/(2\beta+1)}$, which is beyond previous techniques (when the Lipschitz map is involved). Note that the factor $\tfrac{1}{2\beta-1}$ also appears in Dudley’s integral for the supremum of a Gaussian process indexed by the Sobolev ellipsoid, a classical example of the looseness of Dudley’s bound [39], where the optimal dependence is order $\tfrac{1}{\sqrt{2\beta-1}}$.
In [arXiv:2508.18682v2], we further present lifting techniques for tail probability bounds, and application to sparse regression, recovering results of Raskutti et al. [35] for $\ell_{q}$ balls, and further extending the rates to weak $\ell_{q}$-balls. The results for weak $\ell_{q}$-balls appears to be beyond prior techniques.
Further related work: Connections between ERM and upper bounds of stochastic processes, especially characterizing the minimax rate as the solution to a fixed point equation involving the localized Gaussian complexity, can be traced back to [41]. In the random design setting where the symmetrization argument applies, the localized Rademacher complexity was developed by [31, 4, 19]. An alternative approach for bounding the ERM risk using anti-concentration (small ball probability) rather than concentration was developed in [21, 34], which can be used in cases beyond sub-Gaussian noise. On the other hand, for convex regression classes with Gaussian noise, concentration has been used to show that the mean-squared error (MSE) is tightly bounded by the solution to an equation involving a localized Gaussian width [7, 45]. These latter methods crucially rely on convexity of the regression class, hence do not apply to some statistical examples we consider (see Remark 2).
Talagrand’s majorizing measure theorem [37] provides a sharp “geometric” characterization of the supremum of a Gaussian process in term of the $\gamma_{2}$ functional (see definition in (9)). Talagrand also developed several equivalent characterizations of $\gamma_{2}$, including admissible nets, tree-packing and labeled nets; see [39, 42] and the references therein. In particular, the labeled net (called “index” of an atom in the original paper [38, Theorem 2.1]) is closely related to the length of a uniquely decodable code, hence has an information-theoretic interpretation. This connection was clarified and further explored through the Kraft-McMillan inequality by [32, 9]. However, none of these prior works established (1). While general bounds for empirical processes and PAC-Bayes were derived based on $\gamma_{2}$ or its related forms [34, 3, 20], its application to computing minimax rates for specific classes appears to be limited, since $\gamma_{2}$ can often be more difficult to compute than directly evaluating the supremum of a process [38].
A closely related line of research, started by [36] and further developed by other authors including [47, 16, 2, 5, 1], proved information-theoretic generalization bounds in terms of the mutual information between the random index and the process. A basic version (see for example [36]) is
$\mathbb{E}[X_{Z}]\lesssim\sup_{t\in T}\|X_{t}\|_{G}\sqrt{I(Y;Z)}$ (2)
where the exact sub-Gaussian norm $\|\cdot\|_{G}$ will be defined in (4), and $Y:=(X_{t})_{t\in T}$ denotes the stochastic process (in the case of finite $T$, we have $Y\in\mathbb{R}^{|T|}$). While this can be interpreted as a strengthening of a maximal inequality [18], further refinement using the chaining argument, strengthening the Dudley integral, is also possible [2]. This type of results are useful in statistics and machine learning, where $Z$ represents the output of an algorithm which has limited mutual information with the training data, potentially much smaller than the covering number of $T$. Though also concerning a random index framework, their results differ from and are not directly comparable with (1), since their strengthening is in terms of the mutual information between $Z$ and the data, rather than the rate-distortion function of $Z$. Their proofs typically involve information-theoretic arguments such as chain rules and nonnegativity of the relative entropy, but not the lifting technique in the present paper. Nevertheless, we will show in Theorem 2 a result that encompasses both (1) and (2), highlighting the versatility of the lifting technique.
Finally, another direction of extending the theory about supremum of processes, considered recently by [28, 27, 10], generalizes the supremum to a soft-max, also known as free energy, which recovers the supremum as the inverse temperature tends to infinity. It seems an interesting future work to extend the lifting approach to bound the soft-max, and we reported preliminary reports in [27].
Organization: Section 2 recollects standard facts about upper bounds for stochastic processes, its applications to ERM, and relevant information-theoretic tools. Section 3 presents the main results: the rate-distortion integral, its sharpness, and an example of application to ERM over the Sobolev ellipsoid under a 1-Lipschitz map. Section 4-5 prove the upper and lower bounds for the rate-distortion integral. Section 6 presents an equivalent form of the rate-distortion integral which yields a new proof of the majorizing measure theorem, and Section 7 further applies it to the ellipsoid example.
2 Preliminaries
Notation: $\wedge$ and $\vee$ to denote the max and min of two numbers, and $x]_{+}:=x\vee 0$. Given probability measures $P$ and $Q$ on a metric space $(T,d)$, $W_{p}(P,Q):=\sup_{P_{XY}\in\Pi(P,Q)}\mathbb{E}^{\frac{1}{p}}[d(X,Y)^{p}]$ denotes the Wasserstein distance, where $\Pi(P,Q)$ denotes the set of couplings. For brevity, $\Pi(P):=\Pi(P,P)$. Product measures are denoted by $P\times Q$ or $P^{\otimes N}$. The diameter of a set $T$ is ${\rm diam}(T):=\sup_{x,y\in T}d(x,y)$. $\lesssim$ means less than or equal to up to a universal constant, and $\asymp$ means both $\lesssim$ and $\gtrsim$. We denote by $\ell_{2}$ the Hilbert space of square-summable infinite sequences, equipped by the norm $\|\cdot\|_{2}$, abbreviated as $\|\cdot\|$. The empirical distribution of a sequence $x=(x(1),x(2),\dots,x(d))$ is denoted by $\widehat{P}_{x}$. Given a set $T$ in $\mathbb{R}^{d}$ and a matrix $A\in\mathbb{R}^{n\times d}$, define $T-T:=\{t-s\colon t,s\in T\}$.
2.1 Upper bounds for stochastic processes
Classically, Dudley’s integral provides upper bounds on a sub-Gaussian process in terms of the covering numbers. Following [43, Exercise 2.40], we define the sub-gaussian variance of a random variable $X$ as
${\rm var}_{G}(X):=\inf\{\sigma^{2}\colon\mathbb{E}[e^{\lambda(X-\mathbb{E}X)}]\leq e^{\sigma^{2}\lambda^{2}/2}\textrm{ for all $\lambda\in\mathbb{R}$}\},$ (3)
and define the exact sub-Gaussian norm by
$\|X\|_{G}^{2}={\rm var}_{G}(X)+(\mathbb{E}[X])^{2}.$ (4)
This is one of the several equivalent (up to universal constant) definitions of the sub-Gaussian norm; see [43, Proposition 2.6.1]. Let $(X_{t})_{t\in T}$ be a centered (i.e., $\mathbb{E}[X_{t}]=0$) sub-Gaussian process, and define the natural metric
$d(t,t^{\prime}):=\|X_{t}-X_{t^{\prime}}\|_{G},\quad\forall t,t^{\prime}\in T.$ (5)
Then $(T,d)$ is a metric space. Dudley’ integral states that
$\mathbb{E}[\sup_{t\in T}X_{t}]\leq K\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda$ (6)
where $K$ is a universal constant, $\mathsf{N}(T,\lambda)$ denotes the $\lambda$-covering number of $T$, i.e. the minimal number of (open) balls of radius $\lambda$ required to cover $T$, and $\Delta$ denotes the diameter of $T$. While usually called Dudley’s integral, (6) is occasionally attributed to other authors as well; see [15, 33, 39] for more about the history. The Gaussian process is a special case of the sub-Gaussian process where $(X_{t_{1}},X_{t_{2}},\dots,X_{t_{n}})$ follows the Gaussian distribution for any $t_{1},\dots,t_{n}$. For Gaussian processes, a commonly referenced explicit bound is $K<24$ [22].
Recall the following definition of stationary Gaussian processes (see e.g. [42, 39]):
Definition 1.
The Gaussian process $(X_{t})_{t\in T}$ is called stationary if there exists a group $G$ acting on $T$ such that
  1. 1.
    $d(g(t),g(s))=d(t,s)$ for all $t,s\in T$, $g\in G$ (translation invariance).
  2. 2.
    For every $t,s\in T$, there exists $g\in G$ such that $t=g(s)$ (transitivity).
It is well-known that Dudley’s integral is not sharp (i.e., not a two-sided inequality up to universal constants). However, for stationary Gaussian processes, it is known since Fernique [17] that Dudley’s integral is sharp:
$\int_{0}^{\Delta}\sqrt{\ln\mathsf{N}(T,\lambda)}d\lambda\leq K^{\prime}\mathbb{E}[\sup_{t\in T}X_{t}]$ (7)
for a universal constant $K^{\prime}>0$. [33] provides an exposition of the stationary process story and gives a bound of $K^{\prime}\leq 432$. The stationary case is particularly convenient, since the smaller balls in the chaining argument are equivalent in each level, so that the chaining process in the proof of the Dudley integral can be reversed (see [42] for discussions).
For general nonstationary Gaussian processes, the same paper of Fernique [17] conjectured what was later known as the majorizing measure theorem, whose nontrivial lower bound part was proved more than a decade later by Talagrand [37] using a more sophisticated generic chaining argument. Given a probability measure $\mu$ on $T$, define
$I_{\mu}(t):=\int_{0}^{\Delta}\sqrt{\ln\frac{1}{\mu(B(t,\lambda))}}d\lambda.$ (8)
The majorizing measure theorem states that the left side of (6) equals, up to a universal constant factor,
$\gamma_{2}(T):=\inf_{\mu}\sup_{t\in T}I_{\mu}(t).$ (9)
The usual definition of $\gamma_{2}(T)$ in terms of admissible partitions (developed later by Talagrand in the 90s cf. [39]) is equivalent to (9) up to a constant factor. Generally speaking, $\gamma_{2}(T)$ is more difficult to evaluate than Dudley’s integral: $\mu$ needs to be optimized for a quantity depending on $\mu(B(t,\lambda))$ over different scales of $\lambda$, whereas in Dudley’s integral, we can calculate the covering number by optimizing for each scale separately.
Despite this, the majorizing measure theorem is appealing due to its sharpness for the Gaussian process. A nontrivial application of $\gamma_{2}$ is sub-Gaussian comparison: suppose that $(X_{t})$ is a separable sub-Gaussian process, $(Y_{t})$ is a separable Gaussian process, and both processes are zero mean with the same natural metric on $T$. Then
$\mathbb{E}[\sup_{t\in T}X_{t}]\lesssim\mathbb{E}[\sup_{t\in T}Y_{t}].$ (10)
The special case where $(X_{t})$ is also Gaussian is called the Gaussian comparison inequality, usually proved by an interpolation argument which is highly Gaussian-specific [22]. The more general sub-Gaussian case can be immediately seen from the majorizing measure theorem. As noted in [42], this version of sub-Gaussian comparison is a nontrivial consequence of the majorizing measure theorem, for which no other proof is known.
It is also known (see [39]) that the following quantities are equivalent to $\gamma_{2}(T)$ up to constants:
${\rm Fer}(T)$ $:=\sup_{\mu}\int_{T}I_{\mu}(t)d\mu(t);$ (11)
$\delta_{2}(T)$ $:=\sup_{\mu}\inf_{t\in T}I_{\mu}(t).$ (12)
2.2 Empirical risk minimization (ERM)
Consider the mean estimation problem, where $T$ is a closed subset of $\ell_{2}$, and $X_{1},\dots,X_{n}$ are i.i.d. random vectors in $\ell_{2}$ whose mean is an unknown $m\in T$. We assume that $\|X_{1}-m\|_{G}\leq 1$, where following [43], the sub-Gaussian norm of a random vector $Y\in\ell_{2}$ is defined by taking the supremum over one-dimensional projections:
$\|Y\|_{G}:=\sup_{v\colon\|v\|_{2}=1}\|\left<Y,v\right>\|_{G}.$ (13)
A bound on $\|Y\|_{G}$ can be guaranteed, for example, if $Y$ has independent coordinates with uniformly bounded sub-Gaussian norms [43]. For any $t\in\ell_{2}$, define the empirical risk
$\hat{L}(t):=\frac{1}{n}\sum_{i=1}^{n}\|t-X_{i}\|_{2}^{2}.$ (14)
The (population) risk is $L(t):=\mathbb{E}[\hat{L}(t)]$. The empirical risk minimizer (ERM) is defined as
$\hat{m}:=\operatorname*{arg\,min}_{t\in T}\hat{L}(t).$ (15)
We have
$\|\hat{m}-m\|^{2}$ $=L(\hat{m})-L(m)$ (16)
$\leq-\hat{L}(\hat{m})+\hat{L}(m)+L(\hat{m})-L(m)$ (17)
$=\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$ (18)
where $\chi_{z}:=-2\sqrt{n}\left<z,\frac{1}{n}\sum_{i=1}^{n}X_{i}-m\right$>, and $(\frac{1}{2}\chi_{z})_{z\in T-T}$ is a centered sub-Gaussian process, whose natural metric is the $\ell_{2}$ metric.
More generally, the ERM framework also applies to many other problems including regression, functional estimation, matrix completion, shape-restricted regression, and isotonic regression; see [7, 6]. In any case, the inequality (18) plays a key role in upper bounding the mean squared error $\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]$; see e.g. [41, eq (2)]. Since $\mathbb{E}[\hat{L}(m)-L(m)]=0$, the expected generalization error equals $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], which is also connected to (18). If we take the expectation of both sides of (18), and upper bound the right side by the Dudley integral using the covering number of $T$, we generally do not get sharp convergence rates of the mean squared error. In fact, even for the simplest case of 1-dimensional parameter estimation, such a naive approach yields a $1/\sqrt{n}$ rate rather than the optimal $1/n$ rate.
The traditional approach for improving the bound is a “slicing” argument that yields local complexity bounds; see for example [41, 44, 35] or [42, Theorem 5.32]. Roughly speaking, if $\|\hat{m}-m\|^{2}=r$, then $\chi_{m-\hat{m}}\leq\sup_{z\in B(r)\cap(T-T)}\chi_{z}$, which is bounded by some function $\omega(r)$ with high probability. The function $\omega(r)$ typically grows slower than $r^{2}$, so $r^{2}\lesssim\frac{1}{\sqrt{n}}\omega(r)$ yields an upper bound on $r$. More precisely, a union bound argument is applied to control the generalization error at different scales of $r$ with high probability. While it may seem that the uniform control by the union bound could lose logarithmic factors, this can actually be avoided by carefully scheduling of the error tolerance at each scale.
Local complexity bounds are usually based on Dudley’ integral, which can be loose for general $T$ when Dudley’ integral is loose. The majorizing measure theorem, though sharp, is hard to evaluate, and is rarely used in the analysis of ERM.
2.3 Information-theoretic tools
We recall standard information-theoretic measures in [12], using the natural base of the logarithm. The entropy of a discrete random variable $X\sim\mu$ is $H(X):=\sum_{x}\mu(x)\ln\frac{1}{\mu(x)}$. The relative entropy between $\mu$ and $\nu$ on the same measurable space is $D(\mu\|\nu):=\int\ln\frac{d\mu}{d\nu}d\mu$. Mutual information for $(X,Y)\sim P_{XY}$ is defined as $I(X;Y):=D(P_{XY}\|P_{X}\times P_{Y})$. Conditional versions of these quantities are defined by averaging over the random variable which is conditioned on. For example, $D(P_{\hat{Z}|Z}\|\mu|P_{Z}):=\int D(P_{\hat{Z}|Z=z}\|\mu)dP_{Z}(z)$, and $H(Y|X):=\int H(Y|X=x)dP_{X}(x)$, where $H(Y|X=x):=\sum_{y}P_{Y|X=x}(y)\ln\frac{1}{P_{Y|X=x}(y)}$. The lifting argument relies on the method of types, a standard method in information theory [13] and large deviation analysis [14]. The type of a sequence is defined as its empirical distribution. Since not all distributions can be a type for a given length of the sequence, we need the following:
Definition 2.
Suppose that $\mu$ is a distribution on a finite set $\mathcal{Z}$. We say $\mu$ is rational if there exists integer $N>0$ such that $N\mu$ is integer, i.e., $N\mu(z)\in\mathbb{Z}$ for any $z\in\mathcal{Z}$.
If $N\nu$ is integer, the type class of $\nu$ is defined as the set of sequences with type $\nu$. The following result is standard and can be found in [13, Lemma 2.6]:
Lemma 1.
Let $\mathcal{S}_{N}$ be the set of all possible types of $\mathcal{Z}^{N}$, where $\mathcal{Z}$ is a finite set with cardinality $n:=|\mathcal{Z}|$. Then $|\mathcal{S}_{N}|={n+N-1\choose n-1}\leq(N+1)^{n}$. Furthermore, if $\mu$ and $\nu$ are distributions on $\mathcal{Z}$, $N\nu$ is integer, and $\mathcal{C}$ denotes the type class of $\nu$, then
$(N+1)^{-n}\exp(-ND(\nu\|\mu))\leq\mu^{\otimes N}(\mathcal{C})\leq\exp(-ND(\nu\|\mu)).$ (19)
Lemma 1 implies the following result, which can be found in [13, Chapter 2, Problem 3]:
Lemma 2.
Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $P_{XY}$ is a distribution on $\mathcal{X}\times\mathcal{Y}$, and $NP_{XY}$ is integer. Let $x^{N}$ be a sequence of type $P_{X}$, and let $\mathcal{C}\subseteq\mathcal{Y}^{N}$ be the type class of $P_{Y}$. Define
$\mathcal{C}(x^{N}):=\{y^{N}\colon(x^{N},y^{N})\textrm{ has type $P_{XY}$}\}.$ (20)
Then we have
$(N+1)^{-|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq\frac{|\mathcal{C}(x^{N})|}{|\mathcal{C}|}\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (21)
Note that sequences in $\mathcal{C}(x^{N})$ and $\mathcal{C}$ have the same probability under $P_{Y}^{\otimes N}$. Taking $\mu=\nu=P_{Y}$ in Lemma 1 yields
$\quad(N+1)^{-2|\mathcal{X}||\mathcal{Y}|}\exp(-NI(X;Y))\leq P_{Y}^{\otimes N}(\mathcal{C}(x^{N}))\leq(N+1)^{|\mathcal{Y}|}\exp(-NI(X;Y)).$ (22)
We recall the rate-distortion function in information theory [12], which will be a counterpart of the covering number in Dudley’s integral. Given a probability measure $\mu$ on $T$ and $\sigma>0$, define
$\Pi_{\sigma}(\mu):=\{P_{Z\hat{Z}}\colon P_{Z}=P_{\hat{Z}}=\mu;\,\mathbb{E}[d^{2}(Z,\hat{Z})]\leq\sigma^{2}\}.$ (23)
Definition 3.
Given $\mu$ on a metric space $(T,d)$, define the rate-distortion function
$R_{\mu}(\sigma^{2}):=\inf_{P_{UZ}}I(U;Z),$ (24)
where the infimum is over probability measure $P_{UZ}$ on $T\times T$ satisfying $P_{Z}=\mu$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$. Define $r_{\mu}(\sigma):=R_{\mu}(\sigma^{2})$, and
[…]
Observe that
$r_{\mu}(\sigma)\leq i_{\mu}(\sigma)\leq r_{\mu}(\sigma/2).$ (26)
[…]
For $Z\sim\mu$, define
$\sigma_{\rm m}(\mu):=\inf_{z\in T}\sqrt{\mathbb{E}[d^{2}(Z,t)]},$ (27)
abbreviated as $\sigma_{\rm m}$ if no confusion. Clearly $R_{\mu}(\sigma^{2})=0$ whenever $\sigma>\sigma_{\rm m}$. Therefore $\sigma_{\rm m}$ can be thought of as a soft version of the diameter of $T$, since the in the upper limit of the traditional Dudley integral can be set as the diameter.
3 Main results
3.1 A rate-distortion integral
Our first main result concerns $\mathbb{E}[X_{Z}]$, where $Z$ is a random index that may depend on the realization of the process, which is a natural consequence of our information-theoretic technique. We first clarify some measure-theoretic aspects regarding the random index formulation. First, if $T=\{t_{0},t_{1},\dots\}$ is countable, we assume that for each $t\in T$, $X_{t}(\omega)$ is a measurable function of $\omega\in\Omega$ (where $\Omega$ denotes the sample space). Suppose that $Z(\omega)\in\{t_{0},t_{1},\dots\}$ is also a measurable function of $\omega\in\Omega$. Then $X_{Z(\omega)}(\omega)$ is a measurable function of $\omega$, because for any Borel set $B\subseteq\mathbb{R}$,
$\{\omega\colon X_{Z(\omega)}(\omega)\in B\}=\bigcup_{i=0}^{\infty}\{\omega\colon Z(\omega)=t_{i}\}\cap\{\omega\colon X_{t_{i}}(\omega)\in B\}$ (28)
which is measurable.
If $T$ is uncountable, measurability is not so immediate since the union in (28) is uncountable. The standard approach is to assume separability, i.e., there is a dense countable subset $T_{0}=\{t_{0},t_{1},\dots\}\subset T$, and $X_{t}$ is continuous in $t\in T$ almost surely, so that $X_{t}$ for $t\notin T_{0}$ can be defined by passing a limit. Continuity here is with respect to the topology compatible with the metric $d$. We note that separability and continuity ensure that $X_{Z(\omega)}(\omega)$ is measurable, as long as $Z(\omega)$ is measurable, which should suffice for most statistical applications. Indeed, we can construct random variable $Z_{k}:=t_{i}$ where
$i=\operatorname*{arg\,min}_{j\in\{0,1,\dots,k\}}d(Z(\omega),t_{j})$ (29)
with ties broken arbitrarily. Assuming measurability of $Z(\omega)$, we have measurability of $d(Z(\omega),t_{j})$ for each $j$, so that $Z_{k}$, and hence $X_{Z_{k}}$, is measurable. Density of $T_{0}$ ensures $\lim_{k\to\infty}Z_{k}=Z$ almost surely, and continuity implies that $X_{Z}=\lim_{k\to\infty}X_{Z_{k}}$ is measurable, being a sequential limit of measurable functions.
Definition 4.
Given a stochastic process $(X_{t}(\omega))_{t\in T}$ and distribution $\mu$ on $T$, define the complexity
$\mathsf{w}(\mu):=\sup\mathbb{E}[X_{Z}]$ (30)
where the supremum is over all $Z=Z(\omega)$ satisfying $P_{Z}=\mu$.
Clearly $\mathsf{w}(\mu)$ is analogous to the notion of Gaussian complexity of a set.
Theorem 1.
If $(X_{t})_{t\in T}$ is a centered sub-Gaussian process on a countable set $T$, and $\mu$ is a distribution on $T$ for which $\sigma_{\rm m}:=\sigma_{\rm m}(\mu)<\infty$, then
$\mathsf{w}(\mu)\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)}d\sigma\leq 4K\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}d\sigma$ (31)
where $K$ is the constant in (6).
Theorem 1 is proved in Section 4 for finite $T$ and extended to countable $T$ in the appendix. It will be shown the appendix that if the right side of (31) is finite, then $\mathbb{E}[|X_{Z}|]<\infty$, so no need to worry about absolute integrability on the left side of (31).
Another type of information-theoretic improvements of traditional generalization bounds, discussed around (2), have recently been studied in the literature. While this type of results are not directly comparable to Theorem 2, we can use the lifting approach to obtain a strengthening of Theorem 2 that encompass both improvements. For simplicity, assume that $|T|<\infty$, and we have the following improvement of Theorem 1:
Theorem 2.
Suppose that $(X_{t})_{t\in T}$ is a centered sub-Gaussian process, and let $Y:=(X_{t})\in\mathbb{R}^{|T|}$. Let $Z$ be a random variable on $T$ arbitrarily correlated with $Y$, and set $\mu:=P_{Z}$. Let $K$ be the constant in (6). Then
$\mathbb{E}[X_{Z}]\leq 2K\int_{0}^{\sigma_{\rm m}}\sqrt{i_{\mu}(\sigma)\wedge I(Y;Z)}d\sigma.$ (32)
The proof of Theorem 2 is given in Section 4. Since $\sigma_{\rm m}\lesssim\sup_{t\in T}\|X_{t}\|_{G}$, it is easy to see that Theorem 2 implies (2).
While Theorem 1 does not cover the case of uncountable $T$, it is sufficient for our statistical applications due to the following observation:
Proposition 1.
If $(X_{t})_{t\in T}$ is a separable, continuous process, then for $Z\sim\mu$,
$\mathbb{E}[X_{Z}]\leq 2K\lim_{\epsilon\downarrow 0}\sup_{\mu^{\prime}\colon W_{\infty}(\mu,\mu^{\prime})\leq\epsilon}\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu^{\prime}}(\sigma^{2})}d\sigma.$ (33)
The proof of Proposition 1 is in the appendix. Naturally, we would like to show that the right side of (33) reduces to the right side of (31) by dominated convergence, but this requires showing that $\sqrt{R_{\mu^{\prime}}(\sigma^{2})}$ can be dominated by an absolutely integrable function, which is challenging in the regime of vanishing $\sigma$ where the rate-distortion function may grow fast. While it is possible to impose on certain growth condition for dominated convergence to apply, we will not dive into those complications. Instead, in many statistical applications, we will see that $R_{\mu^{\prime}}(\sigma^{2})$ can be controlled using the second moments of $\mu$, so the right side of (33) can be directly evaluated using the continuity of the bound in the second moment, without the need to establish the equivalence of the right sides of (33) and (31).
3.2 Sharpness for Gaussian processes
While the traditional Dudley integral is not sharp, somewhat surprisingly, its information-theoretic counterpart, Theorem 1, is sharp:
Theorem 3.
Suppose that $(X_{t})_{t\in T}$ is a centered Gaussian process, $T$ is a finite set, and $\mu$ is an arbitrary distribution on $T$. Then have
$K^{\prime}\mathsf{w}(\mu)\geq\int_{0}^{\infty}\sqrt{i_{\mu}(\sigma)}d\sigma\geq\int_{0}^{\sigma_{\rm m}(\mu)}\sqrt{R_{\mu}(\sigma^{2})}d\sigma,$ (34)
where $K^{\prime}$ is any universal constant such that (7) holds.
We present a proof of Theorem 3 in Section 5, by using Fernique’s theorem that Dudley’s integral is sharp for stationary Gaussian processes (even though Theorem 3 does not assume stationary process). In the appendix we also extend the result to the case of countable $T$. Theorem 3 suggests that Theorem 1 may yield sharper results in certain statistical applications where the traditional method based on Dudley’s integral is suboptimal. From Theorem 3 we can derive the classical majorizing measure theorem. Recall $I_{\mu}$ defined in (8).
Corollary 1.
For any Gaussian process (not necessarily stationary), we have
$\mathbb{E}[\sup_{t\in T}X_{t}]\geq\bar{c}\inf_{\mu}\sup_{t\in T}I_{\mu}(t),$ (35)
where $\bar{c}:=\sup_{a>1}\frac{\sqrt{1-a^{-2}}}{2aK^{\prime}+\sqrt{2\pi h(a^{-2})}}$, $h(\cdot)$ is the binary entropy function, and $K^{\prime}$ is any universal constant such that (34) holds.
The proof of Corollary 1 is given in Section 6, which is essentially based on lower bounding the rate-distortion function by the probability of a set using the data processing inequality.
Another application of Theorem 3 is sub-Gaussian comparison theorems, which, in statistical applications, sometimes allow us to reduce the problem to simpler “linear” cases where straightforward Cauchy-Schwarz or Hölder inequality arguments apply.
Corollary 2.
Suppose that $(X_{t}(\omega))_{t\in T}$ is a centered sub-Gaussian process on a countable $T$, with complexity function $\mathsf{w}_{1}$ (Definition 4) and natural metric $d_{1}$. Suppose that $(G_{t}(\omega^{\prime}))_{t\in T}$ is a Gaussian process with complexity function $\mathsf{w}_{2}$ and natural metric $d_{2}$, and $d_{1}$ is dominated by $d_{2}$. Then we have $\mathsf{w}_{1}(\mu)\lesssim\mathsf{w}_{2}(\mu)$ for any $\mu$ on $T$.
The proof immediately follows from Theorem 1-3, since the rate-distortion functions under the two metrics can be directly compared. Corollary 2 can be viewed as a “soft” version of (10) (in fact, it is easy to see that it implies (10)). The Euclidean special case of Corollary 2 is of particular interest and will be used later in our statistical applications:
Corollary 3.
If $T$ is a countable subset of $\mathbb{R}^{N}$ and the covariance matrix $\operatorname*{cov}(Z)=\Sigma$, and the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ (see (5)) is dominated by the Euclidean metric, then for any $P_{Z}$ on $T$, we have $\mathbb{E}[X_{Z}]\lesssim\operatorname*{tr}(\sqrt{\Sigma})$.
Proof.
Consider a Gaussian process $(G_{t})_{t\in T}$, where $G_{t}=\left<G,t\right$> and $G\sim\mathcal{N}(0,I_{N})$. The natural metric for this Gaussian process coincides with the Euclidean metric. Then by Cauchy-Schwarz we have $\mathbb{E}[G_{Z}]=\mathbb{E}[\left<\Sigma^{1/4}G,\Sigma^{-1/4}Z\right>]\leq\operatorname*{tr}(\sqrt{\Sigma})$. The claim then follows from Corollary 2. ∎
Although Corollary 3 looks like a simple Cauchy-Schwarz inequality, it would not be true if sub-Gaussianity is replaced by a second-moment bound. More precisely, “the natural metric on $T$ for a centered sub-Gaussian process $(X_{t})_{t\in T}$ is dominated by the Euclidean metric” cannot be replaced by $(X_{t})$ is a stochastic process satisfying $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$”. Indeed, consider $T$ as a maximal $\epsilon$-packing of the unit ball in $\mathbb{R}^{N}$, where $N=3$, and suppose that $X_{t}$ is an independent zero mean random variable with $\epsilon^{2}$ variance for each $t\in T$. Then it is true that $\mathbb{E}[|X_{t}-X_{t^{\prime}}|^{2}]\lesssim\|t-t^{\prime}\|_{2}^{2}$. However, for fixed $N=3$ and vanishing $\epsilon$, it is possible that $\mathbb{E}[X_{Z}]=\mathbb{E}[\max_{t\in T}X_{t}]\asymp\epsilon\sqrt{|T|}\asymp\epsilon^{1-N/2}$ diverges, for certain heavy-tailed $X_{t}$. On the other hand, since $Z$ is in the unit ball, $\operatorname*{tr}(\sqrt{\Sigma})$ remains bounded from above as $\epsilon$ vanishes, confirming that this is a counterexample.
3.3 Application to ERM over the Sobolev ellipsoid
We use the rate-distortion integral to upper bound the error in ERM over a Sobolev ellipsoid under a Lipschitz transformation, A general ellipsoid in $\ell_{2}$ is defined as
$\mathcal{E}:=\left\{t\colon\sum_{i=1}^{\infty}\frac{t_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$ (36)
where $a_{1},a_{2},\dots$ is a decreasing sequence of nonnegative numbers. A Sobolev ellipsoid is an ellipsoid with $a_{i}=i^{-\beta}$ for some $\beta>0$. By considering $t(1),t(2),\dots$ as the Fourier coefficients of a regression function from the $\beta$-Sobolev space, we see that the Sobolev ellipsoid can be identified with a nonparametric function class [40]. Sobolev ellipsoids therefore play a prominent role in nonparametric statistics [41, 44, 40]. The precise problem of ERM over a Sobolev ellipsoid was previously studied in [45].
The ellipsoid is recurring example in [39], in particular an example showing that Dudley’s integral is loose. In Section 7, we review the supremum of a Gaussian process indexed by $\mathcal{E}$, and show that the rate-distortion integral yields a sharp bound.
Consider ERM in the setting of Section 2.2, where $T=\phi(\mathcal{E})$ is the image of the Sobolev ellipsoid (for $\beta>1/2$, where $1/2$ is the critical exponent for ERM) under a 1-Lipschitz map $\phi\colon\ell_{2}\to\ell_{2}$. In practice, $\phi$ may represent a nonlinear neural network or feature map. If $\phi$ is the identity map, the Cauchy-Schwarz inequality suffices for the error upper bound (Remark 2), while the rate-distortion integral machinery is essential for general $\phi$, analogous to the role of the majorizing measure theorem for sub-Gaussian comparison (10). Furthermore, we assume that $Y$ is a random vector in $\ell_{2}$ satisfying $\|Y\|_{G}\leq 1$. Under these assumptions, we have the following main result:
Theorem 4.
There exists a function $c(\beta)>0$ satisfying $\lim_{\beta\downarrow\frac{1}{2}}\sqrt{2\beta-1}c(\beta)<\infty$, such that $\mathbb{E}[\|m-\hat{m}\|_{2}^{2}]\leq c(\beta)n^{-\frac{2\beta}{2\beta+1}}$.
We present a proof of Theorem 4 in Section 7.3.
Remark 1.
Theorem 4 shows that ERM achieves mean squared error of order $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$ with high probability. Furthermore, from the proof of Theorem 4 we can also see that the expected generalization error, $\mathbb{E}[\frac{1}{\sqrt{n}}\chi_{m-\hat{m}}$], also converges at the rate $\frac{1}{\sqrt{2\beta-1}}n^{-\frac{2\beta}{2\beta+1}}$. For comparison, the traditional technique of [41] based on Dudley’s integral will only prove an upper bound at the rate $\frac{1}{2\beta-1}n^{-\frac{2\beta}{2\beta+1}}$ (see Remark 4), which has the optimal exponent of $n$, but a worse preconstant.
Remark 2.
For convex regression classes with Gaussian noise, [7] showed that the mean-squared error (MSE) of ERM can be characterized by a fixed-point equation involving a localized Gaussian width. For ellipsoids, this localized Gaussian width can then be controlled directly via the Cauchy–Schwarz inequality [45], thereby avoiding the slicing and Dudley-integral machinery and mitigating the looseness of Dudley’s bound. However, the Cauchy-Schwarz step of [45] exploits linear structure, and cannot be applied to the setting of Theorem 4 when the map $\phi$ is nonlinear Lipschitz. This limitation is expected to be artifacts of the proof technique of [7, 45]: Lipschitz maps do not increase the supremum of the relevant stochastic process (a consequence of the majorizing measure theorem; see (10)). Our work thus introduces a new proof technique that overcomes this limitation.
4 Proof of the rate-distortion integral: finite case
In this section we prove Theorem 1 and Theorem 2 in the case of finite index set; extension beyond the finite case follows by a limiting argument, and is deferred to the appendix. Recall the following basic property of the sub-Gaussian distribution, whose proof can be found in [43, Exercise 2.40]:
Proposition 2.
Suppose that $X_{1},X_{2},\dots,X_{n}$ are zero-mean, independent random variables, and $\|X_{i}\|_{G}\leq K_{i}$, $i=1,2,\dots,n$. Then $\|X_{1}+X_{2}+\dots+X_{n}\|_{G}\leq\sqrt{K_{1}^{2}+K_{2}^{2}+\dots+K_{n}^{2}}$.
Let $n:=|T|$. Without loss of generality, we can identify $T$ with the standard basis vectors $e_{1},e_{2},\dots,e_{n}$ of $\mathbb{R}^{n}$, and define $Y_{i}:=X_{e_{i}}$ for each $i\in\{1,\dots,n\}$. Then the process can be represented by inner product:
$X_{z}=\left<Y,z\right>,\quad\forall z\in T.$ (37)
It is not necessary to adopt the inner product representation for the proof of Theorem 1, but it helps making the notations clear. Now let $Y^{N}\sim P_{Y}^{\otimes N}$. Recall the definitions related to types in Section 2.3. For any $\mu$ on $T$, define
$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $N\mu$ is integer}\}$ (38)
which is nonempty if $\mu$ is rational. For $N\in\mathcal{I}$, define the type class
$\mathcal{C}_{\mu}$ $:=\{z^{N}\in T^{N}\colon\widehat{P}_{z^{N}}=\mu\}.$ (39)
The following two lemmas show that a coupling can be approximated by a matching:
Lemma 3.
Let $\mu$ be a rational distribution on $\mathbb{R}^{n}$, and let $N\in\mathcal{I}$. For any $y^{N}\in\mathbb{R}^{nN}$, we have
$\frac{1}{\sqrt{N}}\min_{s^{N}\in\mathcal{C}_{\mu}}\|y^{N}-s^{N}\|_{2}=W_{2}(\widehat{P}_{y^{N}},\mu).$ (40)
Lemma 4.
Suppose that $\mathcal{X}$ and $\mathcal{Y}$ are finite sets, $N>0$ is an integer, and $P_{XY}$ is a distribution with the property that both $NP_{X}$ and $NP_{Y}$ are integer. Then there exists $Q_{XY}$ satisfying $Q_{X}=P_{X}$, $Q_{Y}=P_{Y}$, $\max_{x,y}|Q_{XY}(x,y)-P_{XY}(x,y)|\leq\frac{1}{N}$, and that $NQ_{XY}$ is integer.
Lemma 3 has no approximation error, whereas Lemma 4 shows approximation on the level of the distribution (not just the min value) but has a $1/N$ approximation error. For our purpose, it is in fact possible to use only Lemma 4, since the $1/N$ gap vanishes in the limit. The proofs of both lemmas can be found in the appendix.
By Proposition 2, $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a sub-Gaussian process with respect to the following metric on $T^{N}$:
$d_{N}(z^{N},\hat{z}^{N}):=\sqrt{\sum_{i=1}^{N}d^{2}(z_{i},\hat{z}_{i})}.$ (41)
The following observation is crucial:
Lemma 5.
For any rational $\mu$, we have
$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>]=\mathsf{w}(\mu),$ (42)
where $\lim_{\mathcal{I}\ni N\to\infty}$ denotes the limit along the subsequence in $\mathcal{I}$.
Proof.
By Lemma 3, we see that for any $N\in\mathcal{I}$,
$\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<y^{N},z^{N}\right$> $=\frac{\mathbb{E}_{\widehat{P}_{y^{N}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(\widehat{P}_{y^{N}},\mu)}{2}.$ (43)
Define $M_{N}:=\sup_{P_{Y^{\prime}}\colon W_{2}(P_{Y},P_{Y^{\prime}})\leq N^{-\frac{1}{2n}}}\frac{\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]+\mathbb{E}_{\mu}[\|Z\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)}{2}$. Since $W_{2}$ is a metric (satisfying the triangle inequality), we see that $\mathbb{E}_{P_{Y^{\prime}}}[\|Y\|_{2}^{2}]-W_{2}^{2}(P_{Y^{\prime}},\mu)$ is continuous in $P_{Y^{\prime}}$ with respect to $W_{2}$, and hence
$\lim_{\mathcal{I}\ni N\to\infty}M_{N}$ $=\sup_{P_{YZ}\in\Pi(P_{Y},\mu)}\mathbb{E}[\left<Y,Z\right$>] (44)
$=\mathsf{w}(\mu).$ (45)
Define
$\mathcal{D}$ $:=\{y^{N}\in\mathbb{R}^{nN}\colon W_{2}(P_{Y},\widehat{P}_{y^{N}})\leq N^{-\frac{1}{2n}}\}.$ (46)
By the Wasserstein law of large numbers (cf. [8]), we have $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[Y^{N}\in\mathcal{D}]=1$. Therefore,
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]$ $=\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\sup_{P_{YZ}\in\Pi(\widehat{P}_{Y^{N}},\mu)}\mathbb{E}[\left<Y,Z\right>]1_{Y^{N}\in\mathcal{D}}]$ (47)
$\leq\lim_{\mathcal{I}\ni N\to\infty}M_{N}\mathbb{E}[1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu),$ (48)
where the first equality follows from Lemma 3. Similarly, if the $\sup$ in the definition of $M_{N}$ is replaced by $\inf$, then (45) still holds, but the inequality sign in (48) will be reversed, and the $\limsup$ in (47) will be replaced by $\liminf$. Hence we have
$\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}}]=\mathsf{w}(\mu).$ (49)
It remains to show that $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\mathbb{E}[\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>1_{Y^{N}\in\mathcal{D}^{c}}]=0$, which, by the Cauchy-Schwarz inequality, would follow once we show
$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left(\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\right)^{2}\right]<\infty.$ (50)
To show (50), first note that, due to $|T|<\infty$, we have $\|\frac{1}{N}\left<Y^{N},z^{N}\right>\|_{G}\leq\frac{C_{1}}{\sqrt{N}}$ for each $z^{N}\in\mathcal{C}$, where $C_{1},C_{2},C_{3}>0$ here are some constants independent of $N$. Since $\sqrt{\ln|\mathcal{C}_{\mu}|}\leq C_{2}\sqrt{N}$, by [43, Exercise 2.37] we have $\|\frac{1}{N}\max_{z^{N}\in\mathcal{C}_{\mu}}\left<Y^{N},z^{N}\right>\|_{G}\leq C_{3}\cdot\frac{C_{1}}{\sqrt{N}}\cdot C_{2}\sqrt{N}=C_{1}C_{2}C_{3}$. Then (50) follows since the sub-Gaussian norm dominates the square root of the second moment. ∎
We have the following estimate of the covering number $\mathsf{N}(\mathcal{C},\cdot)$ under the metric $d_{N}$:
Lemma 6.
Let $\mu$ be rational, and $\mathcal{C}:=\mathcal{C}_{\mu}$. For any $\sigma>0$, we have
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\leq i_{\mu}(\sigma);$ (51)
$\liminf_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)$ $\geq i_{\mu}(2\sigma).$ (52)
Proof.
For any $t^{N}\in\mathcal{C}$, we have
$\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))=\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (53)
where $\mathcal{S}_{N}$ denotes set of all $P_{Z\hat{Z}}\in\Pi_{\sigma}(\mu)$ such that $NP_{Z\hat{Z}}$ is integer, and $\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}):=\{u^{N}\in{T}^{N}\colon\textrm{$(t^{N},u^{N})$ has type $P_{Z\hat{Z}}$}\}$. Evidently,
$\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq\sum_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))\leq|\mathcal{S}_{N}|\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))$ (54)
where $\lim_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln|\mathcal{S}_{N}|=0$ by Lemma 1. By (22), we have
$\left|\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C}_{P_{Z\hat{Z}}}(t^{N}))-I(Z;\hat{Z})\right|\leq c_{N}$ (55)
where $c_{N}:=\frac{2|{T}|^{2}}{N}\ln(N+1)$. Next, we show that
$\lim_{\mathcal{I}\ni N\to\infty}\max_{P_{Z\hat{Z}}\in\mathcal{S}_{N}}I(Z;\hat{Z})=i_{\mu}(\sigma).$ (56)
Indeed, the $\leq$ part is obvious. For the $\geq$ part, let $P_{Z\hat{Z}}^{*}$ be the optimal distribution in the definition of $\Pi_{\sigma}(\mu,\mu)$. By Lemma 4, there exists $P^{N}_{Z\hat{Z}}\in\mathcal{S}_{N}$ such that $\max_{z,u}|P_{Z\hat{Z}}^{*}(z,u)-P^{N}_{Z\hat{Z}}(z,u)|\leq\frac{1}{N}$. Therefore $\lim_{\mathcal{I}\ni N\to\infty}I(Z;\hat{Z})=i_{\mu}(\sigma)$ and $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{E}[d^{2}(Z,\hat{Z})]=\sigma^{2}$ where $(Z,\hat{Z})\sim P^{N}_{Z\hat{Z}}$. Since $i_{\mu}(\sigma)$ is continuous with respect to $\sigma$, this establishes the $\geq$ part of (56). Now combining (54), (55) and (56), we obtain
$\left|\frac{1}{N}\ln\frac{1}{\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}-i_{\mu}(\sigma)\right|\leq c_{N}^{\prime}$ (57)
where $c_{N}^{\prime}>0$ is a vanishing sequence possibly depending on $\sigma$ and $\mu$. Let ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ denote the packing number, i.e. the maximal number of points $\mathcal{C}$ separated by distance at least $2\sqrt{N}\sigma$; then $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\leq{\sf P}(\mathcal{C},2\sqrt{N}\sigma)$ (see e.g. [42]), and standard volume estimates yields ${\sf P}(\mathcal{C},2\sqrt{N}\sigma)\leq\frac{1}{\min_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},\sqrt{N}\sigma))}$, proving (51). Moreover, by a volume argument we also have $\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)\cdot\max_{t^{N}\in\mathcal{C}}\mu^{\otimes N}(B(t^{N},2\sqrt{N}\sigma))\geq\mu^{\otimes N}(\mathcal{C})$. By Lemma 1, we have $\lim_{N\to\infty}\frac{1}{N}\ln\mu^{\otimes N}(\mathcal{C})=0$. Then using (57) with $\sigma$ replaced by $2\sigma$, we obtain (52). ∎
Proof of Theorem 1 (for finite $T$).
When $|T|<\infty$, it suffices to consider the case of rational $\mu$, since the general case would then follow by a limiting argument. Applying Dudley’s integral (6) to $\mathcal{C}:=\mathcal{C}_{\mu}$ yields
$\mathbb{E}[\max_{z^{N}\in\mathcal{C}}\left<Y^{N},z^{N}\right>]\leq K\int_{0}^{\Delta_{\rm m}}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\,d\sigma$ (58)
where $\Delta_{\rm m}:=\sup_{N\in\mathcal{I}}\frac{{\rm diam}(\mathcal{C})}{2\sqrt{N}}<\infty$. Now we multiply $1/N$ to (58), take $\mathcal{I}\ni N\to\infty$, and apply Lemma 6 and Lemma 5. We can apply the dominated convergence theorem, since $\frac{1}{N}\sqrt{\ln\mathsf{N}(\mathcal{C},2\sqrt{N}\sigma)}2\sqrt{N}\leq\frac{2}{\sqrt{N}}\sqrt{\ln(n^{N})}=2\sqrt{\ln n}$, which is independent of $N$ and absolutely integrable on $\sigma\in[0,\Delta_{\rm m}]$. Then Theorem 1 for finite $T$ follows from Lemma 6, Lemma 5, and (26). ∎
Proof of Theorem 2.
As in the proof of Theorem 1, we can identify $T=\{e_{1},\dots,e_{n}\}$ as the standard basis vectors, so that $X_{z}=\left<Y,z\right$>. Furthermore, we assume that $|\mathcal{Y}|<\infty$ and $P_{YZ}$ is rational; the general case can then be proved by a limiting argument, since both sides of (2) are continuous with respect to the distribution when $|T|<\infty$. Define
$\mathcal{I}:=\{N\in\mathbb{Z}\colon\textrm{$N>0$, $NP_{YZ}$ is integer}\}.$ (59)
For each $N\in\mathcal{I}$, let $\mathcal{C}_{\mu}$ denote the type class of $\mu$, consisting of all sequences of type $\mu=P_{Z}$. Select
$L:=\exp(NI(Y;Z)+N^{1/3})$ (60)
sequences from $\mathcal{C}_{\mu}$ uniformly at random with replacement, and call this random set $\mathcal{A}$. Let $Y^{N}\sim P_{Y}^{\otimes N}$ (independent of this random set). Then we will show that
$\mathbb{E}[\left<Y,Z\right>]\leq\liminf_{\mathcal{I}\ni N\to\infty}\mathbb{E}[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right$>]. (61)
Indeed, set
$\mathcal{B}$ $:=\{(y^{N})\colon\widehat{P}_{y^{N}}=P_{Y}\};$ (62)
$\mathcal{D}$ $:=\{(y^{N},z^{N})\colon\widehat{P}_{y^{N}z^{N}}=P_{YZ}\},$ (63)
and for any $y^{N}\in\mathcal{B}$, define $\mathcal{D}(y^{N}):=\{z^{N}\colon(y^{N},z^{N})\in\mathcal{D}\}$. Standard results about type estimates (immediate from Lemma 2 and Lemma 1) state that:
$\left|\frac{1}{N}\ln|\mathcal{D}(y^{N})|-H(Z|Y)\right|$ $\leq C_{1}\frac{\ln N}{N},\quad\forall y^{N}\in\mathcal{B},$ (64)
$\left|\frac{1}{N}\ln|\mathcal{C}_{\mu}|-H(Z)\right|$ $\leq C_{1}\frac{\ln N}{N},$ (65)
where $C_{1},C_{2}\dots$ in this proof are positive constants independent of $N$. Let $\pi$ be the projection from $\mathbb{R}^{nN}$ to $\mathcal{B}$ (i.e., map to the closest point in $\mathcal{B}$, with ties broken arbitrarily). Define $\mathcal{E}$ as the event that $\mathcal{D}(\pi(Y^{N}))\cap\mathcal{A}=\emptyset$. Then, (64) implies that
$\mathbb{P}[\mathcal{E}]$ $=\left(1-|\mathcal{D}(y^{N})|/|\mathcal{C}_{\mu}|\right)^{L}$ (66)
$\leq\left(1-\exp(-NI(Y;Z)-2C_{1}\ln N)\right)^{L}$ (67)
$\leq\exp(-C_{2}N^{1/3}),$ (68)
where $y^{N}$ is any sequence in $\mathcal{B}$. From the definitions it is clear that
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>1_{\mathcal{E}^{c}}\geq\mathbb{E}[\left<Y,Z\right>]1_{\mathcal{E}^{c}}$ (69)
for all $N\in\mathcal{I}$ almost surely. By our assumption of $|T|,|\mathcal{Y}|<\infty$, we have $\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\leq\max_{y\in\mathcal{Y},z\in T}|\left<y,z\right>|\leq C_{3}$, therefore $\lim_{\mathcal{I}\ni N\to\infty}\mathbb{P}[\mathcal{E}]=0$ and the Cauchy-Schwarz inequality shows that
$\limsup_{\mathcal{I}\ni N\to\infty}\mathbb{E}\left[\left|\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right|\cdot 1_{\mathcal{E}}\right]=0.$ (70)
Then (69) implies
$\liminf_{\mathcal{I}\ni N\to\infty}\left[\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>\right]\geq\mathbb{E}[\left<Y,Z\right$>]. (71)
Furthermore,
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<Y^{N},z^{N}\right>\geq\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N}),z^{N}\right>-\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> (72)
From Lemma 3, we have
$\max_{z^{N}\in\mathcal{A}}\frac{1}{N}\left<\pi(Y^{N})-Y^{N},z^{N}\right$> $\leq\frac{1}{\sqrt{N}}\|\pi(Y^{N})-Y^{N}\|_{2}\cdot\frac{1}{\sqrt{N}}\max_{z^{N}\in\mathcal{A}}\|z^{N}\|_{2}$ (73)
$\leq C_{4}W_{2}(P_{Y},\widehat{P}_{Y^{N}})\cdot$ (74)
which converges to 0 in expectation [8]. Thus (61) is verified by collecting the bounds above.
It remains to apply the Dudley inequality to the right side of (61). Since $\mathcal{A}$ is contained in the set $\mathcal{C}$ in Lemma 6, the upper bound part of Lemma 6 also applies to covering numbers of $\mathcal{A}$. Furthermore, we have the trivial bound $|\mathcal{A}|\leq L$. Therefore,
$\limsup_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\ln\mathsf{N}(\mathcal{A},2\sqrt{N}\sigma)\leq i_{\mu}(\sigma)\wedge I(Y;Z).$ (75)
The claim then follows by applying dominated convergence (see proof of Theorem 1). ∎
5 Proof of sharpness of the rate-distortion integral
In this section we present a proof of Theorem 3 using sharpness of Dudley’s integral for stationary Gaussian processes. We first consider the case where $n:=|T|<\infty$ and $\mu$ is rational. Recall the representation of $X_{z}$ as an inner product in (37), which embeds $Y,Z$ in $\mathbb{R}^{n}$ where $Y$ is a normal vector, and the definitions of $Y^{N}$, $\mathcal{I}$, and $\mathcal{C}:=\mathcal{C}_{\mu}$ as in (37), (38), (39). Then $(\left<Y^{N},z^{N}\right>)_{z^{N}\in\mathcal{C}}$ is a stationary Gaussian process with respect to the metric (41), according to Definition 1, due to invariance under the transitive group of coordinate permutations acting on $\mathcal{C}$. By Lemma 5 and the sharpness of Dudley’s integral for stationary Gaussian processes (7), we see that
$K^{\prime}{\sf w}(\mu)$ $\geq\liminf_{\mathcal{I}\ni N\to\infty}\frac{1}{N}\int_{0}^{\infty}\sqrt{\ln\mathsf{N}(\mathcal{C},\sqrt{N}\sigma)}\sqrt{N}\,d\sigma$ (76)
$\geq\int_{0}^{\infty}\sqrt{i_{\mu}(\epsilon)}d\epsilon$ (77)
where $\mathcal{I}$ is defined in (38), and the last step follows from Lemma 6 and Fatou’s Lemma. The lower bound in terms of $R_{\mu}$ follows from (26). This establishes the theorem for finite $T$ and rational $\mu$. Next, we can drop the rationality assumption under $|T|<\infty$: It is easy to see that both $\mathsf{w}(\mu)$ and $\int_{0}^{\infty}\sqrt{i_{\mu}(\sigma)}d\sigma$ are continuous with respect to $\mu$, so that the rational case implies the general case by a limiting argument. For extension to countable case see the appendix.
6 Penalization form and new proof of the majorizing measure theorem
In this section, we prove Corollary 1. First, observe the following auxiliary result allowing us to write integrals in a “penalized optimization” form, which will be useful not only for proof of Corollary 3 but also in bounding the errors for empirical risk minimization.
Lemma 7.
Suppose that $y\in(0,\infty)$ is a decreasing function of $x\in(0,\infty)$. Then
$2\int_{0}^{\infty}ydx\leq\int_{0}^{\infty}\min_{x}\{\alpha^{-2}x^{2}+y^{2}\}d\alpha\leq 4\int_{0}^{\infty}ydx$ (78)
Proof.
For any $\alpha\in(0,\infty)$, there is unique $x(\alpha)$ satisfying $\alpha=\frac{x(\alpha)}{y(x(\alpha))}$. Then $x$ and $y$ can be parameterized by $\alpha\in(0,\infty)$, and $\int_{0}^{\infty}(\alpha^{-2}x^{2}+y^{2})d\alpha=\int_{\alpha=0}^{\infty}2y^{2}\cdot\frac{ydx-xdy}{y^{2}}=4\int_{0}^{\infty}ydx$. The claim then follows since for any $\alpha$, $\frac{\min_{x\in(0,\infty)}\{\alpha^{-2}x^{2}+y^{2}(x)\}}{\alpha^{-2}x^{2}(\alpha)+y^{2}(x(\alpha))}\in[1/2,1]$. ∎
A similar idea (though in the form of series summation rather than integral) was previously used in proving equivalent forms of the majorizing measure theorem [38]. Here, we find Lemma 7 very convenient in forming information-theoretic functionals that are exactly convex-concave (Remark 3), and also simplifying evaluations of the rate-distortion integral in applications to ERM.
To prove Corollary 1, it suffices to consider the case of finite $T$, since the general case (where the lattice supremum in (35) is defined as $\sup_{T_{0}\subseteq T,|T_{0}|<\infty}\sup_{t\in T_{0}}$; see [39]) will then follows by taking the limit. Then $T$ can be identified with points in $\mathbb{R}^{|T|}$ equipped with the $\ell_{2}$ metric. For $i:=i_{\mu}$ in Definition 3, Lemma 7 implies
$4\sup_{\mu}\int_{0}^{\infty}\sqrt{i(\epsilon)}d\epsilon\geq\sup_{P_{Z}}\int_{0}^{\infty}\min_{P_{Z\hat{Z}}\in\Pi(P_{Z},P_{Z})}\{\alpha^{-2}\mathbb{E}[\|Z-\hat{Z}\|_{2}^{2}]+I(Z;\hat{Z})\}d\alpha$ (79)
where both supremums are over distributions on $T$ and the minimum is over couplings of $P_{Z}$ and $P_{Z}$. The right side of (79) is lower bounded by
$\sup_{P_{Z}}\inf_{\mu}\int_{0}^{\infty}\min_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|Z-U\|_{2}^{2}]+D(P_{U|Z}\|\mu|P_{Z})\}d\alpha,$ (80)
since equality would be achieved if $\inf_{\mu}$ were restricted to $\mu=P_{Z}$ and $\min_{P_{U|Z}}$ were restricted to $P_{U|Z}$ satisfying $P_{U|Z}P_{Z}\in\Pi(P_{Z},P_{Z})$. Now define, for each $z,\alpha$, $F_{\mu}(z,\alpha):=\min_{\nu}\{\alpha^{-2}\mathbb{E}_{U\sim\nu}[\|z-U\|_{2}^{2}+D(\nu\|\mu)]\}$. Since the relative entropy is jointly convex (i.e., $D(\frac{\nu_{1}+\nu_{2}}{2}\|\frac{\mu_{1}+\mu_{2}}{2})\leq\frac{1}{2}D(\nu_{1}\|\mu_{1})+\frac{1}{2}D(\nu_{2}\|\mu_{2})$), we see that $F_{\mu}(z,\alpha)$ is convex in $\mu$. The integral in (80) equals $\int_{0}^{\infty}\mathbb{E}[F_{\mu}(Z,\alpha)]d\alpha$, which is convex in $\mu$ and linear in $P_{Z}$. Then Sion’s minimax theorem implies that (80) equals
$\quad\inf_{\mu}\sup_{P_{Z}}\int_{0}^{\infty}\mathbb{E}[F_{\mu}(Z,\alpha)]d\alpha=\inf_{\mu}\max_{z\in T}\int_{0}^{\infty}F_{\mu}(z,\alpha)d\alpha.$ (81)
Now fix any $\mu$, $z$, $\alpha$, and let $P_{U}:=\nu$ achieve the minimum in the definition of $F_{\mu}(z,\alpha)$. Define $r=a\sqrt{\mathbb{E}[\|z-U\|^{2}]}$, where $U\sim P_{U}$ and $a>1$ can be optimized later. We obtain, from the data processing inequality [12],
$D(P_{U}\|\mu)\geq d(P_{U}(B(z,r))\|\mu(B(z,r)))$ (82)
where $d(a\|b):=a\ln\frac{a}{b}+(1-a)\ln\frac{1-a}{1-b}$ denotes the binary relative entropy function and $B(z,r)$ denotes the ball centered at $z$ with radius $r$. By the Markov inequality we have $P_{U}(B(z,r))\geq 1-a^{-2}$. Then we have
$d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}$ (83)
because if $\mu(B(z,r))>1-a^{-2}$ then the right side of (83) is negative and the inequality is obviously true; if $\mu(B(z,r))\leq 1-a^{-2}$ then by monotonicity of the binary relative entropy function we have $d(P_{U}(B(z,r))\|\mu(B(z,r)))\geq d(1-a^{-2}\|\mu(B(z,r)))=(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}+a^{-2}\ln\frac{a^{-2}}{1-\mu(B(z,r))}$ and (83) also holds. Now the second integral in (81) is lower bounded by
$\int_{0}^{\infty}\inf_{r>0}\left\{\frac{r^{2}}{a^{2}\alpha^{2}}+\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}\right\}d\alpha$ (84)
which, by Lemma 7, is lower bounded by twice of
$\int_{0}^{{\rm diam}(T)}\sqrt{\left[(1-a^{-2})\ln\frac{1-a^{-2}}{\mu(B(z,r))}-a^{-2}\ln a^{2}\right]_{+}}\frac{dr}{a}$
$\geq\int_{0}^{{\rm diam}(T)}\sqrt{(1-a^{-2})\ln\frac{1}{\mu(B(z,r))}}\frac{dr}{a}-\int_{0}^{{\rm diam}(T)}\sqrt{h(a^{-2})}\frac{dr}{a}$ (85)
$=\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}$ (86)
where we used the basic inequality $\sqrt{[x-y]_{+}}\geq\sqrt{x}-\sqrt{y}$ for $x,y\geq 0$. Thus (77) implies
$2K^{\prime}\mathbb{E}[\sup_{z\in T}X_{z}]\geq\inf_{\mu}\sup_{z}\frac{\sqrt{1-a^{-2}}}{a}I_{\mu}(z)-\frac{{\rm diam}(T)}{a}\sqrt{h(a^{-2})}.$ (87)
It is easy to see, by considering two points in $T$, that $\mathbb{E}[\sup_{z\in T}X_{z}]\geq{\rm diam}(T)\mathbb{E}[G\vee 0]=\sqrt{\frac{1}{2\pi}}{\rm diam}(T)$, where $G\sim\mathcal{N}(0,1)$. We finish the proof of Corollary 1 by
$\left(2K^{\prime}+\frac{\sqrt{2\pi h(a^{-2})}}{a}\right)\mathbb{E}[\sup_{z\in T}X_{z}]\geq\frac{\sqrt{1-a^{-2}}}{a}\inf_{\mu}\sup_{z}I_{\mu}(z).$ (88)
Remark 3.
The minimax inequality step (80)-(81) crucially leverages Lemma 7 and the joint convexity of the relative entropy. Specifically, the integral in (80) is convex in $\mu$ and linear in $P_{Z}$, and this property is not true for the functional $\int_{0}^{\infty}\min_{P_{U|Z}\colon\mathbb{E}[\|Z-U\|_{2}^{2}\leq\epsilon^{2}}\sqrt{D(P_{U|Z}\|\mu|P_{Z})}d\epsilon$ before we applied Lemma 7. Moreover, had we not used the relative entropy approach, then the integral in (80) might be replaced by the functional $\int_{0}^{\infty}\min_{r_{Z}}\{\alpha^{-2}\mathbb{E}[r_{Z}^{2}]+\ln\frac{1}{\mu(B(Z,r_{Z}))}\}d\alpha$ (see the $\theta_{i}(T)$ functional defined around [38, Proposition 4.3]), where $P_{U|Z}$ is replaced by the conditional distribution of $\mu$ on the ball $B(Z,r_{Z})$, $r_{Z}$ being a function of $Z$. Again, this functional is not exactly convex in $\mu$, so the application of the minimax theorem is less immediate. A classical approach that has a similar role as (80)-(81) is the Fernique convexity argument, which establishes equivalences (11)-(12) but loses a constant factor [39].
7 Proof of ERM over the Sobolev ellipsoid
In this section we apply the rate-distortion integral to the ellipsoid, and in particular, prove Theorem 4.
7.1 Bounding the rate-distortion integral by variance
A key step in the analysis for the ellipsoid (and possibly many other examples) is to upper bound the rate-distortion integral using the covariance matrix. One possible approach is to show that, under the constraint $\operatorname*{cov}(Z)=\Sigma$, the Gaussian distribution maximizes the rate-distortion function, i.e.
$R_{P_{Z}}(\sigma^{2})$ $\leq R_{\mathcal{N}(0,\Sigma)}(\sigma^{2})$ (89)
for all $\sigma^{2}>0$. While it is possible to prove (89) using Lieb’s rotation argument for proving Gaussian optimality (see [29, 11] for the analysis of a similar problem called the reverse Brascamp-Lieb inequality), we will not provide the details here. Instead, we will use an alternative, simpler approach, leveraging the penalized version of the rate-distortion integral: from Lemma 7, we have
$\int_{0}^{\infty}\sqrt{R_{P_{Z}}(\sigma^{2})}\,d\sigma\leq\frac{1}{2}\int_{0}^{\infty}\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}d\alpha.$ (90)
This penalized form is more convenient, because $P_{U|Z}$ is optimized over a set independent of $P_{X}$, so that saddle point analysis applies, and because exact tensorization holds, allowing direct extension of 1-dimensional bounds to the vector case. Define
$f(t):=\left\{\begin{array}[]{cc}t/2&0\leq t\leq 1\\ \frac{1+\ln t}{2}&t>1\end{array}\right.$. (93)
We have the following:
Lemma 8.
For any distribution $P_{Z}$ on $\mathbb{R}$, we have
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $\leq f(2\alpha^{-2}\sigma_{\rm m}^{2})$ (94)
for any given $\alpha>0$, where $\sigma_{\rm m}^{2}:=\operatorname*{var}(Z)$.
Proof.
By translation invariance, it suffices to assume that $\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}$. Define the functional $f(P_{Z},P_{U|Z}):=\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)$, which is convex in $P_{U|Z}$ and concave in $P_{Z}$, and we have the saddle point condition
$\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\leq f(P_{Z}^{*},P_{U|Z}^{*})\leq\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})$ (95)
where $P_{Z}^{*}=\mathcal{N}(0,\sigma_{\rm m}^{2})$ and $P_{U|Z}^{*}$ is an additive noise Gaussian channel that achieves the infimum; see “Gaussian saddle point” in [12]. Then we have $\inf_{P_{U|Z}}f(P_{Z}^{*},P_{U|Z})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}f(P_{Z},P_{U|Z}^{*})\geq\sup_{P_{Z}\colon\mathbb{E}[|Z|^{2}]\leq\sigma_{\rm m}^{2}}\inf_{P_{U|Z}}f(P_{Z},P_{U|Z})$. This shows that under $\mathbb{E}[|Z|]\leq\sigma_{\rm m}^{2}$, the right side of (90) is maximized by $P_{Z}=\mathcal{N}(0,\sigma_{\rm m}^{2})$, in which case
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[|U-Z|^{2}]+I(U;Z)\}$ $=\inf_{D>0}\left\{\alpha^{-2}D+\frac{1}{2}[\ln\frac{\sigma_{\rm m}^{2}}{D}]_{+}\right\}$ (96)
$=(\frac{1}{2}+\frac{1}{2}\ln\frac{2\sigma_{\rm m}^{2}}{\alpha^{2}})1\{\alpha<\sqrt{2}\sigma_{\rm m}\}+\alpha^{-2}\sigma_{\rm m}^{2}1\{\alpha\geq\sqrt{2}\sigma_{\rm m}\}$ (97)
as desired. ∎
Next, we extend Lemma 8 to the case of vectors followed by a Lipschitz transformation:
Lemma 9.
Assume that $W\in\ell_{2}$ is a random variable and $\Sigma:=\operatorname*{cov}(W)$. Let $Z=\phi(W)$, where $\phi$ is a 1-Lipschitz function, and suppose that $\mathbb{E}[\|Z\|^{2}]\leq\epsilon$. Then, for any $\alpha\in(0,\infty)$,
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (98)
Proof.
By considering $U=0$, it is easy to see that $\inf_{P_{U|X}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\alpha^{-2}\epsilon$. Thus it remains to show that
$\inf_{P_{U|Z}}\{\alpha^{-2}\mathbb{E}[\|U-Z\|^{2}]+I(U;Z)\}\leq\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii}).$ (99)
It suffices to prove (99) in the special case where the 1-Lipschitz map $\phi$ is the identity, since the left side of (99) only decreases with the application of a 1-Lipschitz $\phi$. Furthermore, the left side of (99) is upper bounded by $\sum_{i=1}^{\infty}\inf_{P_{U_{i}|Z_{i}}}\{\alpha^{-2}\mathbb{E}[|U_{i}-Z_{i}|^{2}]+I(U_{i};Z_{i})\}$, because given $P_{U_{i}|Z_{i}}$ we can define $P_{U|Z}=\prod_{i=1}^{\infty}P_{U_{i}|Z_{i}}$, and then $I(U;Z)\leq\sum_{i=1}^{\infty}I(U_{i};Z_{i})$. Then the result in (99) follows from the one-dimensional case, which we already proved in Lemma 8. ∎
7.2 Warm-up: supremum over an ellipsoid
Let $X_{z}:=\left<G,z\right$>, where $G=(G(1),G(2),\dots)$ has i.i.d. standard normal coordinates. The majorizing measure theorem shows that (see [39, Section 2.13]):
$\mathbb{E}[\sup_{z\in\mathcal{E}}X_{z}]\asymp\sqrt{\sum_{i=0}^{\infty}2^{i}a_{2^{i}}^{2}}\asymp\sqrt{\sum_{i=1}^{\infty}a_{i}^{2}}.$ (100)
On the other hand, a nontrivial calculation [39] shows that the Dudley integral is $\asymp\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}$ hence may not be sharp.
Here we show that sharpness is regained using the rate-distortion integral. Suppose that $Z^{\prime}$ is a random variable taking values in $\mathcal{E}$ and $\Sigma:=\operatorname*{cov}(Z^{\prime})$. We have
$\int_{0}^{\infty}\sqrt{R_{P_{Z^{\prime}}}(\sigma^{2})}\,d\sigma$ $\lesssim\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ (101)
$\leq\|(a_{i})_{i=1}^{\infty}\|_{2}$ (102)
where follows (101) from the (90) and Lemma 9, noting that $\int_{0}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\asymp\sqrt{\Sigma_{ii}}$, and (102) follows from the ellipsoid constraint $\sum_{i=1}^{\infty}\frac{\Sigma_{ii}}{a_{i}^{2}}\leq 1$ and the Cauchy-Schwarz inequality. Then for any random variable $Z\in\mathcal{E}$, we have $\mathbb{E}[X_{Z}]\lesssim\|(a_{i})_{i=1}^{\infty}\|_{2}$, which follows from Proposition 1, by taking $T$ and $\mu^{\prime}$ therein to be $\mathcal{E}$ and $P_{Z^{\prime}}$ in (102), respectively.
Let us remark that the bound can also be obtained using Corollary 3, since $\operatorname*{tr}(\sqrt{\Sigma})\leq\sum_{i=1}^{\infty}\sqrt{\Sigma_{ii}}$ by the Schur concavity of the square root function and the fact that the eigenvalues majorizes the diagonal values. Furthermore, the argument here also shows that the same upper bound holds when $T$ is the image of the ellipsoid $\mathcal{E}$ under a 1-Lipschitz map.
For the case of the Sobolev ellipsoid, we see the supremum in (100) is order $\sqrt{\frac{1}{2\beta-1}}$ as $\beta\downarrow 1/2$. In contrast, the Dudley integral is
$\sum_{i=0}^{\infty}2^{i/2}a_{2^{i}}=\sum_{i=0}^{\infty}2^{-\beta i+i/2},$ (103)
giving a worse upper bound of order $\frac{1}{2\beta-1}$ as $\beta\downarrow 1/2$; see [44, Example 5.12] or [39, Section 2.13].
7.3 Error bound for ERM
In this section we prove Theorem 4. Recall (18), where we defined the centered sub-Gaussian process $(\chi_{z})_{T-T}$, and $T$ is now taken to be a 1-Lipschitz image of the Sobolev ellipsoid. Our proof is based on the following key estimate:
Theorem 5.
If $\mathbb{E}[\|Z\|_{2}^{2}]\leq E$ for some $E>0$, then $\mathbb{E}[\chi_{Z}]\lesssim C_{\beta}E^{\frac{2\beta-1}{4\beta}}$ for some function $C_{\beta}>0$ (defined around (124)) satisfying $\lim_{\beta\downarrow\frac{1}{2}}\sqrt{2\beta-1}C_{\beta}<\infty$.
Remark 4.
In [41, Lemma 5.1], the ERM error for the Sobolev space of functions was bounded by localizing the Dudley integral, where $\beta\geq 1$ is assumed an integer. Suppose that we now extend the approach of [41] to the setting of Theorem 4: The covering number is bounded as
$\ln N(\delta,\mathcal{E})\lesssim\delta^{-\frac{1}{\beta}}$ (104)
where we assume throughout this Remark that $\beta$ is bounded above (i.e. we only focus on the behavior of the bound as $\beta\downarrow 1/2$). [41, Example 2.1] estimated the covering number for fixed $\beta$. For the Sobolev ellipsoid $\mathcal{E}$, [44, Example 5.12] provided a proof of (104), from which we can see that the best preconstant in (104) is the order of a constant independent of $\beta$. For local covering numbers, [45] showed that the estimate in (104) still holds, where the preconstant is a universal constant. We can then calculate the Dudley integral up to the scale of the error $\sqrt{E}$, resulting in a bound of $\int_{0}^{\sqrt{E}}\delta^{-\frac{1}{2\beta}}d\delta\lesssim\frac{1}{2\beta-1}E^{\frac{1}{2}-\frac{1}{4\beta}}$. Using our proof instead, we get $\mathbb{E}[\chi_{Z}]\lesssim\frac{1}{\sqrt{2\beta-1}}E^{\frac{1}{2}-\frac{1}{4\beta}}$, which is almost the same, but with a better preconstant. Thus Dudley’s integral results in $\mathbb{E}[\|\hat{m}-m\|^{2}]\lesssim\frac{1}{2\beta-1}n^{-\frac{2\beta}{2\beta+1}}$, where the preconstant $\frac{1}{2\beta-1}$ is worse than $\frac{1}{\sqrt{2\beta-1}}$ in Theorem 4. This gap is related to the looseness of the Dudley integral: as we see in (103), Dudley’s integral gives a bound of order $\frac{1}{2\beta-1}$ on the supremum over $\mathcal{E}$, rather than the optimal $\sqrt{\frac{1}{2\beta-1}}$.
How does the rate-distortion integral improve the preconstant? A naive approach would be to first show that for a random variable $Z\in 2\mathcal{E}$, we have $R_{P_{Z}}(\delta^{2})\lesssim\delta^{-\frac{1}{\beta}}$, where we assume throughtout this paragraph that $\beta$ is bounded above (i.e. we only focus on the behavior of the bound as $\beta\downarrow 1/2$). This is analogous to (104), and we may proceed by plugging it into the rate-distortion integral. But this will still result in a preconstant $\frac{1}{2\beta-1}$ rather than $\frac{1}{\sqrt{2\beta-1}}$. The reason is that the $P_{Z}$ maximizing $R_{P_{Z}}(\delta^{2})$ changes with $\delta$. Instead, we leverage the penalized version (90) and evaluate the integrand therein for given $\operatorname*{cov}(Z)$, calculate the integral, and optimize $\operatorname*{cov}(Z)$ in the end. The proof is made simple by properties of information-theoretic measures such as chain rules of the entropy.
We first need a few preparations. We have
Lemma 10.
$t\int_{0}^{1/t}f(\frac{1}{u^{2}})du\leq 4f(t)$, for all $t>0$.
Proof.
For $0<u_{0}\leq 1$,
$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{0}^{u_{0}}\frac{1+\ln\frac{1}{u^{2}}}{2}du=\frac{3}{2}u_{0}+u_{0}\ln\frac{1}{u_{0}}\leq\frac{3}{2}u_{0}(1+\ln\frac{1}{u_{0}}).$ (105)
For $u_{0}>1$, $\int_{1}^{u_{0}}f(\frac{1}{u^{2}})du=\int_{1}^{u_{0}}\frac{1}{2u^{2}}du=\frac{1}{2}(1-\frac{1}{u_{0}})$, so that
$\int_{0}^{u_{0}}f(\frac{1}{u^{2}})du=\frac{3}{2}+\frac{1}{2}(1-\frac{1}{u_{0}})=2-\frac{1}{2u_{0}}\leq 2.$ (106)
The result follows by taking $u_{0}=1/t$. ∎
Lemma 11.
Suppose that $\sum_{i=1}^{\infty}\Sigma_{ii}i^{\beta}\leq 4$, where $\beta>1/2$. For any E>0{E}>0, we have
$\int_{0}^{\infty}\alpha^{-2}{E}\wedge\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq C_{\beta}{E}^{\frac{2\beta-1}{4\beta}},$ (107)
where $C_{\beta}$ is a constant defined around (124), which satisfies $\lim_{\beta\downarrow 1/2}\sqrt{2\beta-1}C_{\beta}<\infty$.
Proof.
Set $\alpha_{0}:={E}^{\frac{1+2\beta}{4\beta}}$. We have
$\int_{\alpha_{0}}^{\infty}\alpha^{-2}{E}\wedge d\alpha$ $\leq{E}^{\frac{2\beta-1}{4\beta}}.$ (108)
Moreover, by Lemma 10
$\int_{0}^{\alpha_{0}}\sum_{i=1}^{\infty}f(2\alpha^{-2}\Sigma_{ii})d\alpha$ $=\sum_{i=1}^{\infty}\sqrt{2\Sigma_{ii}}\int_{0}^{\frac{\alpha_{0}}{\sqrt{2\Sigma_{ii}}}}f(\frac{1}{u^{2}})du$ (109)
$\leq\alpha_{0}\sum_{i=1}^{\infty}4f(\frac{\sqrt{2\Sigma_{ii}}}{\alpha_{0}}).$ (110)
It will become clear that $\alpha_{0}$ was chosen so that (108) and (110) are both upper bounded by ${E}^{\frac{2\beta-1}{4\beta}}$. Therefore, it remains to consider the optimization problem:
$M_{1}:=\sup_{(\Sigma_{i})}\sum_{i=1}^{\infty}f(\frac{\sqrt{2\Sigma_{i}}}{\alpha_{0}})$ (111)
under the constraint $\sum_{i=1}^{\infty}\Sigma_{i}i^{2\beta}\leq 4$. Also, consider another optimization problem:
$M_{2}:=\sup_{\Sigma(\cdot)}\int_{0}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ (112)
under the constraint $\int_{0}^{\infty}\Sigma(t)t^{2\beta}\,dt\leq 4$, where $\Sigma(\cdot)$ is a nonnegative continuous function. Clearly $M_{2}\geq M_{1}$, since given $(\Sigma_{ii})$ we can define $\Sigma(t)=\Sigma_{\lceil t\rceil}$.
Since $f$ has continuous, strictly increasing first derivative, we can see that an optimizer $\Sigma(\cdot)$ exists, which is continuous. Since $f(\cdot)$ is increasing, $\Sigma(\cdot)$ must be decreasing. There exists unique $t_{0}>0$ such that $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\geq 1$, $t\in(0,t_{0}]$ and $\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}\leq 1$, $t\in[t_{0},\infty)$. Stationarity condition implies that there exists $\lambda>0$ such that
$\frac{1}{4\Sigma(t)}$ $=\lambda t^{2\beta},\quad t\in(0,t_{0}];$ (113)
$\frac{1}{2\alpha_{0}\sqrt{2\Sigma(t)}}$ $=\lambda t^{2\beta},\quad t\in[t_{0},\infty).$ (114)
Setting $t=t_{0}$ in both equations and canceling $\Sigma(t_{0})$, we obtain $\lambda=\frac{1}{2\alpha_{0}^{2}t_{0}^{2\beta}}$. Also, we can express $\Sigma(t)$ as a function of $t$ from the above, and $\int_{0}^{\infty}\Sigma(t)t^{2\beta}\,dt=4$ must be achieved, implying
$4$ $=\int_{0}^{t_{0}}\frac{1}{4\lambda}\,dt+\int_{t_{0}}^{\infty}\frac{1}{8\alpha_{0}^{2}\lambda^{2}t^{2\beta}}\,dt$ (115)
$=\frac{t_{0}}{4\lambda}+\frac{1}{8(2\beta-1)\alpha_{0}^{2}\lambda^{2}t_{0}^{2\beta-1}}$ (116)
$=\frac{\beta}{2\beta-1}\alpha_{0}^{2}t_{0}^{2\beta+1}.$ (117)
Now that we are able to express $t_{0}$ and $\lambda$ as functions of $\alpha$, we can evaluate the optimal value:
$\int_{0}^{t_{0}}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ $=\int_{0}^{t_{0}}\frac{1+\ln\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}}}{2}\,dt$ (118)
$=\int_{0}^{t_{0}}\frac{1+\beta\ln\frac{t_{0}}{t}}{2}\,dt$ (119)
$=\frac{t_{0}(1+\beta)}{2};$ (120)
$\int_{t_{0}}^{\infty}f(\frac{\sqrt{2\Sigma(t)}}{\alpha_{0}})\,dt$ $=\int_{t_{0}}^{\infty}\frac{\sqrt{2\Sigma(t)}}{2\alpha_{0}}\,dt$ (121)
$=\int_{t_{0}}^{\infty}\frac{t_{0}^{2\beta}}{2t^{2\beta}}\,dt$ (122)
$=\frac{t_{0}}{2(2\beta-1)}.$ (123)
In summary, we have shown that $M_{1}\leq M_{2}\leq\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)t_{0}\leq c_{\beta}\alpha_{0}^{-\frac{2}{1+2\beta}}$, where we defined
$c_{\beta}=\left(\frac{1+\beta}{2}+\frac{1}{2(2\beta-1)}\right)\left(\frac{\beta}{4(2\beta-1)}\right)^{-\frac{1}{1+2\beta}}.$ (124)
Hence
$\int_{0}^{\infty}\alpha^{-2}{E}\wedge\sum_{i=1}^{d}f(2\alpha^{-2}\Sigma_{ii})d\alpha\leq 4\alpha_{0}M_{1}+{E}^{\frac{2\beta-1}{4\beta}}=(4c_{\beta}+1){E}^{\frac{2\beta-1}{4\beta}}.$ (125)
The proof is completed by setting $C_{\beta}=4c_{\beta}+1$. ∎
Proof of Theorem 5.
We will apply Proposition 1 with $\mu:=P_{Z}$, and let $\mu^{\prime}$ be an arbitrary distribution supported on TTT-T and satisfying $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$. Then there exists $Z^{\prime}\sim\mu^{\prime}$ such that $\|Z^{\prime}-Z\|_{2}\leq\epsilon$ almost surely, and hence $\mathbb{E}[\|Z^{\prime}\|_{2}^{2}]\leq E+\epsilon^{2}+2\sqrt{E}\epsilon$. From Lemma 9 and Lemma 11 we have
$\int_{0}^{\infty}\inf_{P_{U|Z^{\prime}}}\{\alpha^{-2}\mathbb{E}[\|U-Z^{\prime}\|^{2}]+I(U;Z^{\prime})\}d\alpha\leq C_{\beta}(E+\epsilon^{2}+2\sqrt{E}\epsilon)^{\frac{2\beta-1}{4\beta}}.$ (126)
Then Proposition 1 and (90) shows that
$\mathbb{E}[\chi_{Z}]$ $\lesssim C_{\beta}\lim_{\epsilon\to 0}(E+\epsilon^{2}+2\sqrt{E}\epsilon)^{\frac{2\beta-1}{4\beta}}=C_{\beta}E^{\frac{2\beta-1}{4\beta}}.$ (127)
Proof of Theorem 4.
Taking the expectation of (18) and applying Theorem 5, we obtain
$\mathbb{E}[\|\hat{m}-m\|_{2}^{2}]\lesssim\frac{1}{\sqrt{n}}C_{\beta}\mathbb{E}^{\frac{2\beta-1}{4\beta}}[\|\hat{m}-m\|_{2}^{2}]$ (128)
which implies the claimed result. ∎
8 Acknowledgment
The author would like to thank professor Ramon van Handel for feedback on an early version of the manuscript, and for encouraging the pursuit of a new proof of the majorizing measure theorem based on the rate-distortion integral. The author also thanks professors Garvesh Raskutti, Philippe Rigollet and Yuting Wei for discussion and feedback on the regression examples. This research was supported in part by NSF Grant DMS-2515510.
Appendix A From matching to coupling
A.1 Proof of Lemma 3
The $\geq$ part is obvious from the definition. For the $\leq$ part, fix an arbitrary $\bar{s}^{N}\in\mathbb{R}^{nN}$ satisfying $\widehat{P}_{\bar{s}^{N}}=\mu$. Denote the supports of $\widehat{P}_{y^{N}}$ and $\mu$ by $\mathcal{A}$ and $\mathcal{B}$, respectively. Consider the following function
$F(M)$ $=\frac{1}{N}\sum_{1\leq i,j\leq N}\|y_{i}-\bar{s}_{j}\|_{2}^{2}M_{i,j}$ (129)
$=\frac{1}{N}\sum_{y\in\mathcal{A},\,s\in\mathcal{B}}\|y-s\|_{2}^{2}\sum_{i\colon y_{i}=y}\sum_{j\colon\bar{s}_{j}=s}M_{i,j}$ (130)
on $\mathcal{S}$, which is defined as the set of doubly stochastic matrices $M\in\mathbb{R}^{N\times N}$. We can see from (130) that $\inf_{M\in\mathcal{S}}F(M)\leq W_{2}^{2}(\widehat{P}_{y^{N}},\mu)$. Since $F$ is linear, by the Birkhoff-von Neumann theorem, there exists a permutation matrix $M$ achieving $\inf_{M\in\mathcal{S}}F(M)$. This permutation sends $\bar{s}^{N}$ to an $s^{N}$, achieving $\inf_{M\in\mathcal{S}}F(M)=\frac{1}{N}\|y^{N}-s^{N}\|_{2}$. This proves the $\leq$ part.
A.2 Proof of Lemma 4
The proof follows immediately from the following observation, by taking $A:=[NP_{XY}(x,y)]_{x\in\mathcal{X},y\in\mathcal{Y}}$ and $B:=[NQ_{XY}(x,y)]_{x\in\mathcal{X},y\in\mathcal{Y}}$.
Lemma 12.
Let $A\in\mathbb{R}_{\geq 0}^{m\times n}$ have integer row sums and integer column sums. Then there exists $B\in\mathbb{Z}_{\geq 0}^{m\times n}$ with the same row and column sums as $A$ and
$\max_{i,j}\,|B_{ij}-A_{ij}|\;\leq\;1.$
Proof.
Write $F:=A-\lfloor A\rfloor$, so $F_{ij}\in[0,1)$ for all $i,j$. Let $r_{i}:=\sum_{j}F_{ij}$ and $c_{j}:=\sum_{i}F_{ij}$. Because the row and column sums of $A$ are integers and $\sum_{j}\lfloor A_{ij}\rfloor,\sum_{i}\lfloor A_{ij}\rfloor$ are integers, we have $r_{i},c_{j}\in\mathbb{Z}$.
Form the bipartite graph $G$ with left vertices $\{1,\dots,m\}$ (rows) and right vertices $\{1,\dots,n\}$ (columns), and put an edge $(i,j)$ iff $F_{ij}\in(0,1)$. If some connected component of $G$ were a tree, it would have a leaf $v$ incident to a unique edge $e$ of weight $F_{e}\in(0,1)$, forcing the corresponding $r_{i}$ or $c_{j}$ to equal $F_{e}\notin\mathbb{Z}$, a contradiction. Hence every component of $G$ contains a cycle.
Fix a simple cycle $C$ in $G$. Along the edges of $C$, alternately add $\varepsilon$ and $\varepsilon$ to the corresponding entries of $F$ (leaving all other entries unchanged), where
$\varepsilon\;:=\;\min\{\,F_{ij},\,1-F_{ij}:(i,j)\in C\,\}.$
This preserves all row and column sums, since at each vertex on $C$ the two incident edges change by $\varepsilon$ and $\varepsilon$. Moreover, all modified entries remain in $[0,1]$, and at least one becomes 0 or $1$. Remove edges for which $F_{ij}$ becomes 0 or 1. Iterating this operation, which terminates in finitely many steps, we obtain a modified matrix $B$ with the desired properties. ∎
Appendix B Extension of the rate-distortion integral beyond finite $T$
In the main text, we proved the rate-distortion integral for finite $T$. In this section, we extend the proof beyond the finite case by a limiting argument. We also prove Proposition 1.
B.1 Absolute integrability of $|X_{Z}|$
We first prove an auxiliary result about absolute integrability of $|X_{Z}|$, which will be used later to apply dominated convergence.
Lemma 13.
Suppose that $(X_{t})_{t\in T}$ is a zero-mean sub-Gaussian process on a countable set $T$. Let $Z\sim\mu$ be a random variable on $T$, for which $\sigma_{\rm m}:=\sigma_{\rm m}(\mu)<\infty$.
  • Consider $\bar{T}:=T\times\{-1,1\}$, and a process $(\bar{X}_{\bar{t}})_{\bar{t}\in\bar{T}}$, where $\bar{X}_{\bar{t}}:=sX_{t}$ for $\bar{t}=(t,s)$. Let $\bar{d}$ be the natural metric for this sub-Gaussian process on $\bar{T}$, and $\bar{Z}:=(Z,S)$ be a random variable on $\bar{T}$, satisfying $P_{Z}=\mu$. Define $\bar{\sigma}_{\rm m}^{2}:=\inf_{\bar{t}\in\bar{T}}\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{t})]$, and let $\bar{R}_{P_{\bar{Z}}}(\cdot)$ be the rate-distortion function for $\bar{Z}$ on $(\bar{T},\bar{d})$. Then
    $\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}}}(\sigma^{2})}d\sigma\leq\int_{0}^{\sigma_{\rm m}}\sqrt{R_{Z}(\sigma^{2})}\,d\sigma+\sqrt{2\ln 2}\sigma_{\rm m}.$ (131)
  • If $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}\,d\sigma<\infty$, then $\mathbb{E}[|X_{Z}|]<\infty$.
Lemma 13 may be compared with the classical observation that for symmetric processes, $\mathbb{E}[\sup_{t\in T}|X_{t}|]<\infty$ is equivalent to $\mathbb{E}[\sup_{t\in T}X_{t}]<\infty$, both equivalent to boundedness of the process in the Gaussian case; see [37, p2], [39, Exercise 2.2.2.] and the references therein.
Proof.
We first prove (131). The natural metric for the new process satisfies $\bar{d}((t,s),(t^{\prime},s))=d(t,t^{\prime})$, for any $s\in\{-1,+1\}$ and $t,t^{\prime}\in T$. Now suppose that $P_{\bar{Z}}=P_{Z,S}$ is an arbitrary distribution on $\bar{T}$. We can assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. By the triangle inequality we can show that $\bar{d}(\bar{Z},(t_{0},+1))\leq d(Z,t_{0})+\|2X_{t_{0}}\|_{\Psi_{2}}=d(Z,t_{0})$, implying that $\bar{\sigma}_{\rm m}^{2}\leq\mathbb{E}[\bar{d}^{2}(\bar{Z},(t_{0},+1))]<\sigma_{\rm m}^{2}$. In addition, suppose that $U$ is any random variable such that $R_{P_{Z}}(\sigma^{2})=I(U;Z)$ and $\mathbb{E}[d^{2}(U,Z)]\leq\sigma^{2}$ for some $\sigma^{2}$. If we set $\bar{U}=(U,S)$, we have $\mathbb{E}[\bar{d}^{2}(\bar{Z},\bar{U})]=\mathbb{E}[d^{2}(Z,U)]\leq\sigma^{2}$, and moreover, the rate-distortion function on $\bar{T}$ satisfies $\bar{R}_{P_{\bar{Z}}}(\sigma^{2})\leq I(\bar{U};\bar{Z})=I(U,S;Z,S)=I(U;Z)+I(S;Z|U)+H(S|Z)\leq R(\sigma^{2})+2\ln 2$. Then (131) follows by the sub-additivity of the square root function.
It remains to prove the second claim. Suppose that $T$ is enumerated as $t_{0},t_{1},\dots$, assuming without loss of generality $X_{t_{0}}\equiv 0$, $\sigma_{\rm m}=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$, and write $T_{k}:=\{t_{0},\dots,t_{k}\}$. Given $Z=Z(\omega)$, define $Z_{k}=Z$ if $Z\in T_{k}$ and $Z_{k}=t_{0}$ otherwise. We can show that
$\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\leq\sqrt{R_{P_{Z}}(\sigma^{2})+2\ln 2}\leq\sqrt{R_{P_{Z}}(\sigma^{2})}+\sqrt{2\ln 2}.$ (132)
Indeed, if $P_{ZU}$ is an optimal distribution in the definition of $R_{P_{Z}}(\sigma^{2})$, then setting $E=1\{Z\notin T_{k}\}$ and $U_{k}=U$ if $E=0$, $U_{k}=t_{0}$ otherwise, we have
$d(U_{k},Z_{k})\leq d(U,Z),\quad\textrm{a.s.}$ (133)
and
$I(U_{k};Z_{k})$ $\leq I(U_{k},E;Z_{k})$ (134)
$=I(E;Z_{k})+I(U_{k};Z_{k}|E)$ (135)
$\leq H(E)+I(U_{k};Z_{k}|E)$ (136)
$\leq H(E)+I(U;Z|E)$ (137)
$\leq H(E)+I(UE;Z)$ (138)
$\leq 2H(E)+I(U;Z),$ (139)
establishing (132). Now applying the result of Theorem 1 in the finite case, we have
$\mathbb{E}[|X_{Z_{k}}|]$ $=\mathbb{E}[{\rm sign}(X_{Z_{k}})X_{Z_{k}}]$ (140)
$\lesssim\int_{0}^{\bar{\sigma}_{\rm m}}\sqrt{\bar{R}_{P_{\bar{Z}_{k}}}(\sigma^{2})}\,d\sigma$ (141)
$\lesssim\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\,d\sigma+\sqrt{2\ln 2}\sigma_{\rm m},$ (142)
where we defined $\bar{Z}_{k}:=(Z_{k},{\rm sign}(X_{Z_{k}}))$, and (142) follows from (131) by replacing the arbitrary $P_{Z}$ with $P_{Z_{k}}$ and noting that $\sigma_{\rm m}(P_{Z_{k}})\leq\sigma_{\rm m}(P_{Z})=:\sigma_{\rm m}$. Thus (132) and (142) imply that $\lim_{k\to\infty}\mathbb{E}[|X_{Z_{k}}|]<\infty$. The proof is completed since $\lim_{k\to\infty}|X_{Z_{k}}|=|X_{Z}|$ almost surely, and Fatou’s lemma shows $\mathbb{E}[|X_{Z}|]\leq\lim_{k\to\infty}\mathbb{E}[|X_{Z_{k}}|]$. ∎
B.2 Extension of Theorem 1 beyond finite $T$
We extend the result of Theorem 1 from the finite case to the countable case. We can assume that $\int_{0}^{\sigma_{\rm m}}\sqrt{R_{\mu}(\sigma^{2})}\,d\sigma<\infty$, since otherwise there is nothing to prove. Again, assume without loss of generality that there is $t_{0}\in T$ satisfying $X_{t_{0}}\equiv 0$ and $\sigma_{\rm m}:=\sigma_{\rm m}(P_{Z})=\sqrt{\mathbb{E}[d^{2}(t_{0},Z)]}$. Construct $Z_{k}$ and define $E$ as in the proof of Lemma 13. Lemma 13 implies $\mathbb{E}[|X_{Z}|]<\infty$, and since $|X_{Z_{k}}|\leq|X_{Z}|\vee|X_{t_{0}}|$, dominated convergence implies
$\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]=\mathbb{E}[\lim_{k\to\infty}X_{Z_{k}}]=\mathbb{E}[X_{Z}].$ (143)
From the definition of $Z_{k}$, we see that $\sigma_{\rm m}(P_{Z_{k}})\leq\sigma_{\rm m}(P_{Z})=:\sigma_{\rm m}$. Hence $R_{P_{Z_{k}}}(\sigma^{2})=0$ for all $\sigma\geq\sigma_{\rm m}$. Then from the estimate in (132), we see that $\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}$ is dominated by the absolutely integrable function $(\sqrt{R_{P_{Z}}(\sigma^{2})}+\sqrt{2\ln 2})1_{\sigma\leq\sigma_{\rm m}}$. Furthermore, since $P(E)$ converges 0 as $k\to\infty$, (139) implies that $\lim_{k\to\infty}R_{P_{Z_{k}}}(\sigma^{2})=R_{P_{Z}}(\sigma^{2})=0$ for any $\sigma>0$. Therefore dominated convergence shows
$\lim_{k\to\infty}\int_{0}^{+\infty}\sqrt{R_{P_{Z_{k}}}(\sigma^{2})}\,d\sigma=\int_{0}^{\sigma_{\rm m}}\sqrt{R_{P_{Z}}(\sigma^{2})}\,d\sigma.$ (144)
Then the claim of the lemma follows by (143), (144) and applying the result of Theorem 1 in the finite case (already proved in the main text).
B.3 Proof of Proposition 1
Let $T_{0}$ be the countable subset in the definition of separability. For any $\epsilon>0$, let $Z_{\epsilon}\in T_{0}$ be a random variable such that $|X_{Z}-X_{Z_{\epsilon}}|\leq\epsilon$ and $d(Z,Z_{\epsilon})\leq\epsilon$, which is possible by the separability and continuity assumption. Then $\mathbb{E}[X_{Z}]\leq\mathbb{E}[X_{Z_{\epsilon}}]+\epsilon$, and $W_{\infty}(\mu,\mu^{\prime})\leq\epsilon$, if $\mu^{\prime}$ is the distribution of $Z_{\epsilon}$. Therefore the claim follows by the result for the countable case and taking $\epsilon\downarrow 0$.
B.4 Extension of Theorem 3 beyond finite $T$
We now use a limiting argument to extend the result to the case of $\mu$ supported on a countable $T$. Assume that $T=\{t_{0},t_{1},\dots\}$. Let $T_{k}:=\{t_{0},t_{1},\dots,t_{k}\}$, $\mu_{k}(t_{i}):=\frac{\mu(t_{i})}{\mu(T_{k})}1\{i\leq k\}$, and $\tilde{\mu}_{k}(t_{i}):=\frac{\mu(t_{i})}{1-\mu(T_{k})}1\{i>k\}$. If $Z_{k}(\omega)$ maximizes $\mathbb{E}[X_{Z_{k}}]$ subject to $P_{Z_{k}}=\mu_{k}$, we can construct a Bernoulli random variable $E\sim{\rm Bern}(\mu(T_{k}))$ independent of $\omega$, and let $Z=Z_{k}$ when $E=1$, and sample $Z\sim\tilde{\mu}_{k}$ independently of $\omega$ when $E=0$. Then $Z\sim\mu$, and
$\mathbb{E}[X_{Z}]=\mu(T_{k})\mathbb{E}[X_{Z_{k}}].$ (145)
Moreover, if $P_{U_{k}|Z_{k}}$ achieves the infimum in the definition of $R_{\mu_{k}}(\sigma^{2})$, we can construct $P_{U|Z}$ by setting $P_{U|Z}(|z)=P_{U_{k}|Z_{k}}(|z)$ for $z\in T_{k}$ and $U=t_{0}$ if $z\notin T_{k}$. Assume without loss of generality that $\mathbb{E}_{\mu}[d^{2}(t_{0},Z)]=\sigma_{\rm m}^{2}$. Then $\mathbb{E}[d^{2}(U,Z)]=\sigma^{2}\mu(T_{k})+\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]$. Since $\lim_{k\to\infty}\mathbb{P}[T_{k}^{c}]=0$ and $\mathbb{E}[d^{2}(t_{0},Z)]<\infty$, a basic result from measure theory (which can be proved by dominated convergence) states that $\lim_{k\to\infty}\mathbb{E}[d^{2}(t_{0},Z)1_{Z\in T_{k}^{c}}]=0$. In particular, for any given $\sigma>0$, there exists $k(\sigma)>0$ such that for all $k>k(\sigma)$, this construction of $P_{U|Z}$ ensures that $\mathbb{E}[d^{2}(U,Z)]<(1+a)\sigma^{2}$, where $a>0$ is an arbitrary constant. Moreover, if $E_{k}:=1\{Z\in T_{k}\}$, we have $I(U;Z)\leq I(U;Z|E)+H(E)=R_{\mu_{k}}(\sigma^{2})+h(\mu(T_{k}))$, where $h(\cdot)$ is the binary entropy function. Taking $k\to\infty$ we have $\mu(T_{k})\to 1$ and obtain
$\lim_{k\to\infty}R_{\mu_{k}}(\sigma^{2})\geq R_{\mu}((1+a)\sigma^{2})$ (146)
for any $\sigma>0$. Then using (145) we obtain
$K^{\prime}\mathbb{E}[X_{Z}]\geq K^{\prime}\lim_{k\to\infty}\mathbb{E}[X_{Z_{k}}]\geq\lim_{k\to\infty}\int_{0}^{\infty}\sqrt{R_{\mu_{k}}(\sigma^{2})}\,d\sigma\geq\int_{0}^{\infty}\sqrt{R_{\mu}((1+a)\sigma^{2})}d\sigma$ (147)
by Fatou’s lemma. Finally, taking $a\downarrow 0$ shows that the countable case holds with the same constant as the finite case.